### v1

# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

Abstract
Understanding human mobility through Point-of-Interest (POI) recommendation is increasingly important for applications such as urban planning, personalized services, and generative agent simulation. However, progress in this field is hindered by two key challenges: the over-reliance on older datasets from 2012-2013 and the lack of reproducible, city-level check-in datasets that reflect diverse global regions. To address these gaps, we present Massive-STEPS (Massive Semantic Trajectories for Understanding POI Check-ins), a large-scale, publicly available benchmark dataset built upon the Semantic Trails dataset and enriched with semantic POI metadata. Massive-STEPS spans 12 geographically and culturally diverse cities and features more recent (2017-2018) and longer-duration (24 months) check-in data than prior datasets. We benchmarked a wide range of POI recommendation models on Massive-STEPS using both supervised and zero-shot approaches, and evaluated their performance across multiple urban contexts. By releasing Massive-STEPS, we aim to facilitate reproducible and equitable research in human mobility and POI recommendation. The dataset and benchmarking code are available at: https://github.com/cruiseresearchgroup/Massive-STEPS.
1 Introduction
Importance of Human Mobility Data and Modeling
Human mobility data and modeling are essential for understanding how individuals interact with and move through physical spaces. This understanding enables a wide range of applications, including urban planning yuan2025worldmove , travel service recommendations feng-etal-2025-agentmove , improved commercial advertising strategies 10.1145/3477495.3531983 , and Point-of-Interest (POI) recommendation 10.1145/3369822 ; 10.1145/3626772.3657840 ; zhang2025survey . Recently, human mobility data has become even more crucial with the increasing use of large language models (LLM) agents to simulate human-like behavior and routines zhou2024large ; jiawei2024large . However, while simulated human mobility data are starting to gain popularity 10.1145/3624475 ; 10.1145/3394486.3412862 , they may not accurately reflect real-world human behavior SALIM2020106964 , highlighting the value of evaluating on real-world data. These advancements are enabled by and large with the rise of Location-based Social Networks (LBSNs), which generate vast amounts of spatio-temporal data through user check-ins  zhang2025survey ; 10.1145/3626772.3657840 . This rich data source has allowed the development of POI recommendation systems that leverage users’ historical visiting behaviors to suggest relevant locations. Such systems enhance user engagement through personalization and provide commercial value to both users and businesses by aligning recommendations with individual preferences and available services zhang2025survey ; 10.1145/3369822 .
Literature Gaps
Our paper addresses three critical gaps in POI recommendation research and datasets, as detailed in Section 2.1. First, as shown in Fig. 1, the field remains dominated by studies focused on just two cities, New York and Tokyo, based on the Foursquare dataset curated in yang2014modeling . This dataset, collected in 2012-2013, raises concerns about its temporal quality, as many POIs may no longer exist and user behavior may have changed yeow2021point . While some recent studies have expanded to other cities 10.1007/s11280-024-01279-y ; merinov2024positive ; feng-etal-2025-agentmove , they often rely on the Global-scale Check-in Dataset (GSCD) yang2015nationtelescope ; yang2016participatory , which, despite its large coverage, is also from 2012-2013 and contains nearly 50% erroneous entries monti2018semantic .
Second, most existing studies are difficult to reproduce, either due to the lack of clearly defined geographic boundaries or the unavailability of the datasets themselves, hindering fair comparison and replication. Finally, we join recent efforts yuan2025worldmove in advocating for the inclusion of low-resource and underrepresented cities in evaluation. Expanding beyond well-studied urban centers is essential for building more generalizable and universally applicable POI recommendation models. Table 1 summarizes these limitations in terms of geographic coverage, temporal span, and reproducibility.
Table 1: Comparison of check-in datasets commonly used for POI recommendation tasks. GSCD  yang2014modeling ; yang2016participatory and Semantic Trails monti2018semantic are global datasets not grouped into individual cities, whereas others perform city-level grouping. Replicable indicates whether city boundaries are clearly defined or can be reliably reconstructed.
Dataset Scale xxCompletenessxx Usability
#cities Years #months POI Attributes Replicable Open-source
GSCD yang2014modeling ; yang2016participatory Varies 2012-2013 17    Coordinates, Category N/A
Semantic Trails monti2018semantic Varies 2012-2013, 2017-2018 24    Category N/A
NYC and Tokyo yang2014modeling 2 2012-2013 11    Coordinates, Category
Gowalla-CA cho2011friendship ; 10.1145/2484028.2484030 1 2009-2010 21    Coordinates, Category
AgentMove feng-etal-2025-agentmove 12 2012-2013 17    Coordinates, Category
Massive-STEPS 12 2012-2013, 2017-2018 24 Coordinates, Category, Name, Address
Massive-STEPS Dataset
In this paper, we introduced the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) Dataset, derived from the Semantic Trails dataset (STD) monti2018semantic . Massive-STEPS includes high-quality check-ins from 2012-2013 and 2017-2018, providing more modern and updated POI check-in data. This supports longitudinal POI recommendation studies and addresses the limitations of older datasets commonly used in prior studies. The dataset covers 12 diverse cities across multiple regions, including East, West, and Southeast Asia, North and South America, Australia, the Middle East, and Europe. Notably, we placed a deliberate emphasis on under-explored regions by including cities such as Jakarta, Kuwait City, and Petaling Jaya, filling a key gap in POI recommendation research that has largely focused on major urban centers. We further enriched STD by aligning it with Foursquare’s Open Source Places dataset, incorporating metadata such as POI coordinates, POI names, and addresses—details unavailable in the original STD.
Benchmark Tasks
To demonstrate the utility of this dataset, we conducted an extensive benchmark on two tasks: (1) supervised POI recommendation and (2) zero-shot POI recommendation. Our benchmark covers a wide range of models, including traditional approaches, deep learning-based models, and more recent LLM-based methods. The goal of POI recommendation task is to predict a set of POIs that a user is likely to visit based on their current check-in trajectory and historical behavior. This reflects real-world applications such as personalized POI recommendations in location-based services. In addition, the scale of our dataset allows us to examine how urban features influence POI recommendation accuracy. Building on prior hypotheses, we propose a new insight: cities with more evenly distributed POI categories tend to be harder to model, as the absence of a dominant POI category makes user behavior less predictable.
Contribution
This paper introduces the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset, addressing gaps in existing POI recommendation research. Current POI check-in datasets are often only from 2012-2013, skewed to a few cities, and lack semantic metadata, hindering the development of robust and globally applicable models. While datasets like GSCD and STD offer broad geographic coverage, they either suffer from an older timespan, contain erroneous data, or have missing information. Massive-STEPS overcomes these issues by providing high-quality check-ins from 2012-2013 and 2017-2018, improving temporal quality for longitudinal POI recommendation studies. The dataset spans 12 diverse cities across multiple regions, with a focus on low-resource cities overlooked in previous research. Additionally, Massive-STEPS is enriched with metadata through alignment with Foursquare’s Open Source Places, providing crucial details such as POI geographical coordinates, POI names, and addresses. We also conducted an extensive benchmark on both supervised and zero-shot POI recommendation tasks, evaluating a wide range of models, including traditional methods, deep learning approaches, and recent LLM-based techniques. We further analyzed which urban features affect POI recommendation accuracy and found that cities with no dominant POI category tend to be harder to predict. By releasing this dataset and benchmark code publicly, we facilitate open and reproducible research, enabling future advancements in POI recommendation studies.
2 Related Works
Figure 1: Distribution of POI recommendation studies modeled on specific cities, filtered from Table IV in zhang2025survey . We identified and counted studies that explicitly mentioned city names, revealing the skewness of existing research, which is saturated around New York and Tokyo. In addition, we include the distribution of studies by LBSN platform, showing that Foursquare is by far the most commonly used source of check-in data. The list of identified studies is shown in Table 5.
2.1 Existing Datasets
A survey conducted by zhang2025survey outlines the broad landscape of POI recommendation research, covering a wide range of models and architectures used in prior studies. While it offers a high-level overview of the datasets used, it lacks a dedicated discussion or evaluation of POI datasets. We address this gap by analyzing commonly used datasets in the field and positioning our dataset within this context.
LBSN Check-in Data Sources
Building on the tabular summary provided in zhang2025survey , which offers a representative overview of the broader POI recommendation literature, we investigated which datasets are most commonly used in prior studies. From their original table (Table IV), we filtered entries pertaining specifically to POI and next POI recommendation tasks and identified (1) the most frequently used LBSN check-in data sources and (2) the most commonly studied cities. As shown in Fig. 1, Foursquare remains the dominant source of LBSN data in existing studies, appearing in almost 50% of the surveyed works. While several variants of Foursquare datasets have been employed, the most widely used are the NYC and Tokyo Dataset yang2014modeling (often abbreviated as FSQ-NYC and FSQ-TKY) and the Global-scale Check-in Dataset (GSCD) yang2015nationtelescope ; yang2016participatory , curated by the same authors. Other LBSN sources occasionally used include Gowalla cho2011friendship , Brightkite cho2011friendship , and Weeplaces 10.14778/3115404.3115407 .
Saturated to Two Cities and Old Timespan
Due to the widespread use of FSQ-NYC and FSQ-TKY yang2014modeling , the majority of POI recommendation studies are disproportionately focused on these two cities, as illustrated in Fig. 1. While there is nothing inherently problematic about studying NYC and Tokyo, there has been growing interest in expanding research to a broader range of cities, particularly those that are underexplored or considered low-resource yuan2025worldmove , as cultural and regional differences influence collective mobility behaviors. For instance, in some urban contexts, residents tend to commute to central business districts in the morning, whereas in others, nightlife activities such as visiting bars after work are more common yang2015nationtelescope . Ensuring diverse geographic coverage is increasingly important, especially as LLMs are adopted for POI recommendation tasks. LLMs are known to exhibit geographical biases, often underperforming in regions with lower socioeconomic conditions manvi2024large . It remains an open question whether LLM-based POI recommendation models can effectively generalize across diverse urban environments.
In addition, because many studies rely on the FSQ-NYC and FSQ-TKY, they are often constrained to the timespan it covers: check-in data from 2012 to 2013. However, POI data is inherently dynamic: venues may have closed, relocated, or changed in category over time. yeow2021point underscores the importance of validating the temporal quality of POI datasets by recording whether and when a venue’s information has been updated to reflect real-world changes. This is particularly critical, as recommender systems should avoid suggesting POIs that no longer exist or have undergone substantial changes (e.g., a former bookstore converted into a coworking space) and behave dynamically over longitudinal periods Yabe2024 . Moreover, behavioral patterns captured over a decade ago may no longer align with modern user preferences and routines. For example, the opening of a new train station may significantly shift commuting patterns and the popularity of surrounding POIs.
Low Data Quality: Erroneous Entries
More recently, researchers have begun leveraging the broader Global-scale Check-in Dataset (GSCD) yang2015nationtelescope ; yang2016participatory , which spans 415 cities across 77 countries. Despite its wider geographic coverage, GSCD is temporally limited to the same 2012-2013 period as FSQ-NYC and FSQ-TKY, and thus suffers from similar issues of temporal quality. More critically, monti2018semantic demonstrated that GSCD suffers from significant data quality issues, with over 14 million check-ins (about 44%) of the dataset flagged as erroneous due to anomalous user behavior. These include (1) repeated check-ins at the same venue, (2) check-ins occurring within implausibly short time intervals (less than one minute), and (3) transitions between venues that would require travel speeds exceeding Mach 1, which are physically unreasonable.
To address these limitations, monti2018semantic introduced the Semantic Trails Dataset (STD), which applies systematic filtering procedures to enhance data quality. STD comprises two subsets: a cleaned version of GSCD covering 2012-2013 (STD 2013), and a newer collection of check-ins from 2017-2018 (STD 2018), sourced from Foursquare Swarm. STD 2018 also spans a wider range of cities, making it valuable for capturing globally distributed user behavior, in contrast to GSCD’s focus on densely populated urban centers. Both subsets follow the same rigorous filtering criteria, resulting in a higher-quality check-in dataset for POI recommendation tasks. Given these improvements, we adopted STD as the source for our check-in dataset.
Poor Reproducibility
Another persistent challenge in POI recommendation research is the lack of reproducibility in dataset preprocessing. While some recent studies utilize datasets like GSCD to cover a wide range of cities, they often omit important details needed for replicating their data filtering processes. For example, AgentMove feng-etal-2025-agentmove conducted city-level filtering based on a minimum distance between each trajectory and the city center, but they did not specify how the city center was defined or what distance threshold was used. Generally, city filtering can be done in two ways: (1) by selecting points within a fixed radius from a central coordinate, or (2) by applying precise administrative boundaries, such as those from OpenStreetMap. The radius-based approach is less reliable, as it assumes a circular city shape and requires a subjective definition of the city’s center. In contrast, the boundary-based method is more robust and reproducible, and is the approach we adopted in this work. Without standardized, transparent filtering criteria and publicly available datasets, it remains difficult to ensure fair comparisons across POI recommendation methods.
2.2 Understanding Urban Features and POI Recommendation
POI recommendation studies that evaluate models across multiple city-level datasets often include analyses to assess how well their methods generalize across different urban contexts. It is well understood that POI recommendation accuracy metrics (e.g., Acc@k, NDCG@k) can vary substantially between cities and can be interpreted as a proxy for how easy or difficult a city is to model. The assumption is that higher performance reflects more predictable or structured mobility patterns. This viewpoint is consistent with prior work highlighting the role of cultural and urban-specific factors in shaping mobility behaviors yang2015nationtelescope ; sun2024city .
Several studies have proposed hypotheses connecting specific urban features to modeling difficulty. For example, GETNext yang2022getnext hypothesized that cities with fewer check-ins and higher spatial sparsity of POIs are harder to model. STHGCN 10.1145/3539618.3591770 suggested that a larger number of user trajectories improves predictive accuracy by providing richer collaborative signals, whose architecture is designed to leverage. LLM4POI 10.1145/3626772.3657840 proposed that cities with a greater variety of POI categories are easier to model due to LLMs’ contextual reasoning capabilities, whereas cities covering a broader geographic area tend to be more difficult to model. In the zero-shot POI recommendation setting, AgentMove  feng-etal-2025-agentmove reported two key findings: (1) geospatial biases inherent in LLMs can hinder prediction quality across cities, and (2) LLMs are influenced by city-specific mobility patterns.
Building on these insights, we used Massive-STEPS to explore how urban features affect POI recommendations. Its diverse set of 12 cities allows for a comprehensive analysis across different cultural and urban contexts. We analyzed the correlation between urban features and model accuracy, and based on the results, proposed a new hypothesis that contrasts previous findings in the literature.
3 Dataset
Table 2: Summary statistics of the 12 Massive-STEPS subsets, including the number of users, trajectories, POI locations, total check-ins, and train, validation, and test sample counts. For comparison, we also include statistics from existing Foursquare-based yang2014modeling and Gowalla-based cho2011friendship datasets. Due to variations in dataset preprocessing across studies, we report the version used in 10.1145/3539618.3591770 .
City Users Trajectories POIs Check-ins #train #val #test
NYC and Tokyo Check-in Dataset yang2014modeling
New York 1,048 14,130 4,981 103,941 72,206 1,400 1,347
Tokyo 2,282 65,499 7,833 405,000 274,597 6,868 7,038
Gowalla cho2011friendship ; 10.1145/2484028.2484030
California 3,957 45,123 9,690 238,369 154,253 3,529 2,780
Massive-STEPS
Beijing 56 573 1,127 1,470 400 58 115
Istanbul 23,700 216,411 53,812 544,471 151,487 21,641 43,283
Jakarta 8,336 137,396 76,116 412,100 96,176 13,740 27,480
Kuwait City 9,628 91,658 17,180 232,706 64,160 9,166 18,332
Melbourne 646 7,864 7,699 22,050 5,504 787 1,573
Moscow 3,993 39,485 17,822 105,620 27,639 3,949 7,897
New York 6,929 92,041 49,218 272,368 64,428 9,204 18,409
Petaling Jaya 14,308 180,410 60,158 506,430 126,287 18,041 36,082
São Paulo 5,822 89,689 38,377 256,824 62,782 8,969 17,938
Shanghai 296 3,636 4,462 10,491 2,544 364 728
Sydney 740 10,148 8,986 29,900 7,103 1,015 2,030
Tokyo 764 5,482 4,725 13,839 3,836 549 1,097
3.1 Creation Process
Massive-STEPS is derived from the two subsets of STD monti2018semantic , incorporating check-ins from both the 2013 and 2018 subsets. We utilize two additional components from STD: (1) the cities metadata file, which provides the latitude and longitude of administrative regions (e.g., towns, suburbs) along with their corresponding country codes obtained from GeoNames; and (2) the POI category mapping, which links each Foursquare Category ID to its descriptive name (e.g., "Restaurant"). Based on this metadata, each POI is thus associated with several attributes: Foursquare Place ID, Foursquare Category ID, category name, latitude/longitude of the administrative region, the administrative region name, and the country code. For anonymization purposes and model training compatibility, we apply ordinal encoding to the Place IDs and Category IDs, assigning each a unique integer index.
Figure 2: World map highlighting the cities included in the Massive-STEPS dataset.
3.1.1 Preprocessing
Trajectory Grouping
Most POI recommendation models operate on sequences of check-ins, commonly referred to as trajectories. The model is tasked with predicting the next POIs a user is likely to visit, given the current trajectory. STD conveniently provides pre-grouped trajectories (trails) by applying a time interval-based grouping: for each user, check-ins that occur within a time interval of $\delta_{\tau}=8$ hours are grouped into the same trajectory.
Matching Trajectories to Target Cities
To obtain city-specific datasets, we matched trajectories to the target cities. For each city, we obtain geographic boundaries from OpenStreetMap and retrieve its GeoJSON file via the Overpass API. The GeoJSON file contains a polygon defining the city’s boundary in latitude and longitude. Using this boundary, we filter check-ins by comparing the latitude/longitude of each POI’s administrative region and retain only those that are within the city’s polygon. This ensures that all retained trajectories are spatially grounded within the designated city.
Filtering Short Trajectories and Inactive Users
To ensure data quality, we apply an additional filtering step by removing trajectories with fewer than two check-ins and excluding users with fewer than three trajectories. This prevents the model from learning from overly sparse or irrelevant data.
Train, Validation, and Test Splits
We split trajectories into training, validation, and test sets in a ratio of 7:1:2, following feng-etal-2025-agentmove . We ensure that all users in the test set appear at least once in the training or validation set, a common practice in prior studies yang2022getnext ; 10.1145/3539618.3591770 .
3.1.2 POI Enrichment via Foursquare OS Places
Since the POIs in STD include their corresponding Foursquare Place IDs, we matched them directly with entries in the Foursquare OS Places dataset using these IDs as the key. This one-to-one ID correspondence allows for a straightforward join operation, enriching each POI with additional metadata such as its precise latitude and longitude, name (e.g., of a restaurant or subway station), and address. However, not all POIs in the Foursquare OS Places dataset include the full metadata, particularly those categorized as private residences, which are excluded due to privacy restrictions.
3.2 Description and Addressing Literature Gaps
Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset is a city-level POI check-in dataset derived from Semantic Trails monti2018semantic , comprising user check-in trajectories from 12 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, Moscow, New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. It features anonymized POI check-ins enriched with geographical metadata to support spatiotemporal and sequential modeling tasks. City-level statistics, along with comparisons to existing datasets, are presented in Table 2. Fig. 2 shows a world map highlighting the locations of all cities included in the Massive-STEPS dataset. Table 6 shows the available fields in the dataset and provides an example for each field.
Massive-STEPS offers a more comprehensive and diverse representation of urban mobility compared to typical POI check-in datasets. As shown in Table 2, datasets like FSQ-NYC and FSQ-TKY yang2014modeling contain fewer than 10,000 candidate POI locations. In contrast, cities in Massive-STEPS cover significantly more POIs: Massive-STEPS New York has over 49,000 POIs, while Massive-STEPS Jakarta exceeds 76,000. Massive-STEPS Istanbul, one of the largest subsets, features a large user base of 23,700, offering a broad range of user behaviors. Although some Massive-STEPS subsets are smaller than their FSQ counterparts (e.g., Tokyo), we attribute this to the strict filtering procedures applied by STD to remove erroneous entries, as explained in Section 2.1. This scale introduces additional computational challenges. Models that rely on dense POI-to-POI adjacency matrices, for instance, may require more efficient implementations to avoid excessive memory consumption.
Beyond scale, Massive-STEPS addresses the oversaturation of FSQ-NYC and FSQ-TKY in POI recommendation research. Notably, Massive-STEPS includes low-resource and previously underexplored cities in human mobility studies, such as Petaling Jaya and Kuwait City, both of which are among the cities with the highest number of check-ins from STD. This broader coverage opens new research opportunities for studying location-based behaviors across diverse cultural and geographic contexts. Furthermore, since Massive-STEPS is based on STD, it benefits from the carefully filtered, high-quality check-ins and a longer, more recent timespan. These characteristics make Massive-STEPS a more relevant and reliable resource for modeling human mobility patterns.
Massive-STEPS is designed to be easily extended to other geographical regions. Since the data processing code is open-source and fully reproducible, adding a new city only requires its geographic boundaries from OpenStreetMap. Moreover, Massive-STEPS is scalable to higher levels of geographic granularity, enabling the creation of provincial, state, and country-level POI check-in datasets, which support collective mobility studies at broader geographic scales.
4 Benchmark
[…]
Model Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
FPMC 0.000 0.026 0.029 0.021 0.062 0.059 0.032 0.026 0.030 0.084 0.075 0.176
RNN 0.085 0.077 0.049 0.087 0.059 0.075 0.061 0.064 0.097 0.055 0.080 0.133
[…]
STHGCN 0.453 0.241 0.197 0.225 0.168 0.223 0.146 0.174 0.250 0.193 0.227 0.250
4.1 POI Recommendation
This benchmark covers the common task of POI recommendation, where the goal is to predict where a user will go next based on their previous check-ins. The input to the model is a trajectory, which is a sequence of places the user has visited. The model is expected to suggest a set of $K$ POIs that the user might visit next. This task can also be framed as re-ranking a set of $N$ candidate POIs in the city based on the user’s recent behavior. This task reflects a typical real-world use case in POI recommendation systems and human mobility modeling. It is a supervised task as the model is trained using all available historical trajectories of the user, allowing it to learn personalized patterns of movement over time. Appendix C provides the details on problem formulation, hyperparameters, experimental setups, and the full results with all evaluation metrics.
Dataset Preparation
By design, our dataset is structured to support consistency and reproducibility, making it straightforward to use for supervised POI recommendation tasks. We adopted the predefined trajectories from the original STD, which has grouped sequences of check-ins into trajectories based on a fixed time interval (see Section 3.1.1). Since all input features in our dataset have been numerically encoded to facilitate model training, we simply used them across all experiments.
For this task, models typically use four kinds of features: (1) social features: user ID; (2) spatial features: POI ID, geographic coordinates; (3) temporal features: check-in timestamp; and (4) categorical features: POI category. Since not all POIs have exact geographic coordinates (see Section 3.1.2), we use the geographic coordinates of their administrative region as a proxy for all POIs.
Models
We evaluated three kinds of models: (1) Markov-based methods, (2) classical deep learning models, and (3) Transformer-based graph neural networks:
Evaluation Metrics
We used two popular metrics in POI recommender systems: Acc@k, which checks if the true POI appears in the top-k predicted results, and NDCG@k, which measures the ranking quality of the suggested results.
Results
As shown in Table 3, STHGCN consistently achieves the highest accuracy across all cities, highlighting the effectiveness of Transformer-based GNNs. GETNext often ranks second, performing well in most cities, and older RNN-based models such as LSTPM and DeepMove also remain competitive in a few cities. The best-performing model achieves an average Acc@1 of 22.9%, which is comparable to results from previous studies using datasets of a similar size feng-etal-2025-agentmove .
We investigated which urban features affect POI recommendation accuracy by computing Spearman correlations between various city features and model performance. We found that category entropy, based on Shannon entropy, strongly correlates with recommendation accuracy ($r=-0.736$), as shown in Fig. 6. Namely, cities with more evenly distributed POI categories and no dominant POI category tend to have lower accuracy and are thus harder to predict. This finding is consistent with the results of prior work on other datasets. Further details are provided in Appendix D.
4.2 Zero-shot POI Recommendation
An increasingly popular approach in POI recommendation leverages the zero-shot capabilities of LLMs, allowing models to perform zero-shot POI recommendation without any additional training. This setting challenges LLMs to generalize to any target city and personalize recommendations for any user. The task mirrors the supervised counterpart: given a user’s trajectory, the model predicts the user’s next destination by ranking a set of $K$ POIs. Appendix E provides the details on problem formulation, prompts, experimental setups, and the full results with all evaluation metrics.
Dataset Preparation
Since LLMs are central for this task, trajectories must be transformed into textual prompts 10.1145/3557915.3561026 ; 10.1109/TKDE.2023.3342137 . We adapted the prompts used in feng-etal-2025-agentmove , which implemented the three LLM methods evaluated in our study. Notably, because LLMs can effectively leverage contextual information, not all input features need to be numerically encoded. The evaluated methods utilize only a subset of features: the timestamp, POI category name, and POI ID.
Methods and Models
We evaluated three LLM-based prompting methods:
  • LLM-Mob wang2023would : One of the earliest methods to use LLMs for next POI prediction, prompting LLMs with both historical and current (contextual) trajectories.
  • LLM-ZS beneduce2024large : A simplified version of LLM-Mob that retains the use of historical and contextual trajectories but simplifies its prompt design.
  • LLM-Move feng2024move : Extends previous prompting methods by introducing a RAG-like approach, retrieving nearby POIs as candidates, and ranking them by geographic distance to the user’s most recent visit.
To ensure a more robust evaluation, we tested each method using four different LLMs: one closed-source API (Gemini 2.0 Flash geminiteam2024geminifamilyhighlycapable ) and three open-source models (Qwen 2.5 7B qwen2.5 , Llama 3.1 8B  grattafiori2024llama3herdmodels , and Gemma 2 9B gemmateam2024gemma2improvingopen ). All open-source models are instruction-tuned variants, quantized to INT4 using AWQ lin2024awqactivationawareweightquantization , and served through vLLM kwon2023efficient .
Evaluation Metrics
We used the same metrics as in the supervised setting, Acc@k and NDCG@k, for the same reasons: to assess predictive accuracy and to evaluate the quality of the ranking.
Results
As shown in Table 4, LLM-Move feng2024move performed the best among the three methods, likely due to its well-crafted prompts that supply potential candidate POIs, rather than relying solely on historical or contextual trajectories like LLM-Mob and LLM-ZS. This also supports the use of administrative region coordinates as a proxy when exact POI coordinates are unavailable (e.g., due to privacy issues). Across different LLMs, Gemini 2.0 Flash achieved the highest accuracy across all three prompting strategies, with Qwen 2.5 7B and Gemma 2 9B following closely as strong open-source alternatives. In contrast, we found that Llama 3.1 8B often struggled to follow prompt instructions and frequently produced irrelevant predictions, especially when using the prompts of LLM-Mob and LLM-Move.
Notably, these zero-shot methods outperformed, matched, or came close to the performance of supervised baselines in several cities (e.g., Jakarta, Kuwait City, Moscow), demonstrating their effectiveness even without additional fine-tuning. These results indicate that LLMs can effectively leverage the contextual information included in trajectory prompts to guide their predictions. Since LLMs are pre-trained, they do not require additional training, which reduces both computation time and cost. While inference may require more powerful hardware, it can still be faster overall than training supervised models from scratch.
[…]
Method LLM Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
[…]
Qwen 2.5 7B 0.058 0.035 0.105 0.080 0.030 0.090 0.070 0.030 0.090 0.040 0.035 0.110
Llama 3.1 8B 0.000 0.020 0.055 0.030 0.010 0.030 0.025 0.010 0.030 0.005 0.020 0.005
[…]
Qwen 2.5 7B 0.038 0.040 0.065 0.050 0.040 0.080 0.050 0.045 0.095 0.045 0.045 0.120
Llama 3.1 8B 0.077 0.040 0.045 0.060 0.040 0.080 0.055 0.030 0.030 0.060 0.040 0.110
Gemma 2 9B 0.096 0.045 0.105 0.070 0.050 0.080 0.075 0.065 0.075 0.050 0.045 0.110
[…]
Llama 3.1 8B 0.058 0.015 0.015 0.010 0.040 0.005 0.035 0.040 0.045 0.020 0.055 0.030
[…]
5 Limitations and Future Work
Firstly, Massive-STEPS is derived from the Semantic Trails dataset and thus inherits its biases and potential errors, which may propagate through downstream tasks. Additionally, the dataset is sparse in several cities, which can impact model training quality and limit cross-city generalization. Secondly, Massive-STEPS focuses solely on trajectories and POI metadata, without including user demographic or social information (e.g., age, social connections). This restricts its applicability for personalized or socially-aware POI recommendation tasks. Thirdly, while our benchmarking covers a wide range of models and cities to emphasize replicability and geographic breadth, we did not perform extensive hyperparameter tuning, which may affect the peak performance of the evaluated models.
While the current version of Massive-STEPS already surpasses existing POI check-in datasets in scale and diversity, we envision expanding it further to include even more cities worldwide. Our dataset creation pipeline is designed to be easily extensible, allowing researchers to integrate additional cities with minimal effort. This allows Massive-STEPS to evolve into a truly global POI check-in dataset.
6 Conclusion
In this paper, we presented the Massive-STEPS dataset to address longstanding limitations in POI recommendation research, particularly the reliance on older, geographically saturated, and non-reproducible check-in datasets. Massive-STEPS offers a large-scale, semantically enriched resource spanning 12 cities across diverse global regions and two time periods, supporting both longitudinal and cross-city analyses. The dataset includes rich semantic information such as venue name, address, category, and coordinates. We also provide benchmark results for both supervised and zero-shot POI recommendation methods, illustrating the dataset’s utility across model types and urban contexts. By releasing Massive-STEPS and our evaluation pipeline publicly, we aim to advance open, reproducible, and globally inclusive research in human mobility and POI recommendation systems. We released our dataset under the Apache 2.0 license.
Acknowledgments and Disclosure of Funding
We would like to thank the support of the National Computational Infrastructure and the ARC Center of Excellence for Automated Decision Making and Society (CE200100005). Computational facilities were provided by the School of Computer Science and Engineering at UNSW Sydney through the Wolfpack computational cluster. Additionally, we express our gratitude to the NVIDIA Academic Grant Program for providing access to A100 GPUs on Saturn Cloud. We further thank the Google Cloud Research Credits program for API access to Gemini models.
Appendix A Existing POI Recommendation Datasets
To examine the trend of the usage of POI recommendation datasets, we filtered the comprehensive survey by [68] to extract studies that explicitly mention the cities used in their experiments. The resulting distribution is summarized in Table 5, which shows a strong concentration of studies focused on New York and Tokyo. Additionally, Fig. 1 visualizes the same data, highlighting the uneven distribution of city choices across studies. We also include information on the LBSN platforms used, revealing that Foursquare remains the predominant data source in the field. These findings underscore the need for broader, more inclusive datasets that support evaluation across a wider range of global cities.
Table 5: Overview of POI Recommendation Studies by City and LBSN Platform. This table is adapted from Table IV in the survey by [68] and presents a filtered list of POI recommendation studies that explicitly mention city names and their associated LBSN platforms.
Study Cities LBSN
SSTPMF [5] New York, Tokyo Foursquare, Gowalla
ST-LSTM [69] California, Singapore Brightkite, Foursquare, Gowalla
LSMA [43] New York, San Fransisco, Tokyo Foursquare, Weeplaces
DLAN [46] New York, Tokyo Foursquare
TLR-M [13] New York, Tokyo Foursquare
GETNext [60] New York, Tokyo, California Foursquare, Gowalla
CARAN [17] New York, Tokyo Foursquare, Gowalla
JANICP [70] New York, Tokyo Foursquare, Weeplaces
Li et al. [23] New York, Tokyo Foursquare
AMACF [58] New York, Tokyo Foursquare, Weeplaces
CHA [64] New York, Tokyo Foursquare
HAT [47] Beijing, Shanghai Yelp, others
STAR-HiT [49] New York Foursquare, Gowalla
CAFPR [14] Tokyo, California, Budapest, Melbourne Foursquare
TGAT [19] New York, Tokyo Foursquare
MobGT [50] New York Foursquare, Gowalla
POIBERT [16] Budapest, Delhi, Edinburgh, Glasgow, Osaka, Perth, Toronto Flickr
AutoMTN [30] New York, Tokyo Foursquare
CCDSA [40] New York, Tokyo, San Fransisco Foursquare, Weeplaces
TDGCN [2] Tokyo, California Foursquare, Gowalla, Weeplaces
BayMAN [48] New York Foursquare, Gowalla
ROTAN [11] New York, Tokyo, California Foursquare, Gowalla
STGCN [15] Boston, Chicago, London Gowalla, others
ADQ-GNN [44] New York, Tokyo Foursquare, Gowalla
HS-GAT [67] Boston, Chicago, London Yelp, others
HKGNN [66] New York, Jakarta, Kuala Lumpur, Sao Paulo Foursquare
S2GRec [24] New York, Tokyo Foursquare, Gowalla
GSBPL [39] New York, Tokyo Foursquare, Gowalla
LSPSL [18] New York, Tokyo Foursquare
SCL [3] Florence, Rome, Pisa, Edinburgh, Toronto Flickr
LLM-Move [10] New York, Tokyo Foursquare
LLM4POI [22] New York, Tokyo, California Foursquare, Gowalla
DiffPOI [31] Singapore, New York, Tokyo Foursquare, Gowalla
DSDRec [45] New York, Tokyo Foursquare
Diff-DGMN [72] Istanbul, Jakarta, Sao Paulo, New York, Los Angeles Foursquare
Appendix B Data Visualization
We present several visualizations highlighting Massive-STEPS’ scale and diversity to complement our dataset description.
In Fig. 3, we show the top 10 most frequent POI categories for each city. The distribution reflects the local culture and lifestyle across different urban areas. For example, Beijing and Shanghai have a high number of Chinese restaurants, while Melbourne and Sydney show a strong presence of cafes. In Tokyo, convenience stores and ramen shops dominate. These patterns illustrate the diversity of local culture and user interests.
Fig. 4 plots the distribution of trajectory lengths (i.e., number of check-ins per trajectory). The distribution is long-tailed, with most trajectories being relatively short, similar to the original Semantic Trails dataset. This indicates that users often make only a few check-ins per outing.
Finally, we show the distribution of user activity levels, measured by the number of trajectories per user in Fig. 5. Most users exhibit cold-start behavior, contributing only a small number of trajectories. This highlights the importance of models that are robust to sparse and short user histories.
Table 6: Fields available in the Massive-STEPS dataset, including user, POI, geographic/spatial, and temporal details, along with example data for each field.
Field Description Example
trail_id Numeric identifier of trajectory 2013_2866
user_id Numeric identifier of user 90
venue_id Numeric identifier of POI venue 185
latitude Latitude of POI venue -33.87301862604473
longitude Longitude of POI venue 151.20668402700997
name POI name Sydney Town Hall
address Street address of POI venue 483 George St
venue_category POI category name City Hall
venue_category_id Foursquare Category ID 4bf58dd8d48988d129941735
venue_category_id_code Numeric identifier of POI category 72
venue_city Administrative region name Sydney
venue_city_latitude Latitude of administrative region -33.86785
venue_city_longitude Longitude of administrative region 151.20732
venue_country Country code AU
timestamp Check-in timestamp 2012-04-22 08:20:00
Figure 3: Top 10 most frequent POI categories in each city, highlighting local cultural and urban preferences.
Figure 4: Distribution of trail lengths, showing a long-tailed pattern with most trajectories consisting of a few check-ins.
Figure 5: Distribution of user activity based on the number of trajectories per user, indicating a cold-start-heavy dataset.
Appendix C POI Recommendation: Task Details
We adopt the conventional problem formulation used in prior POI recommendation studies  [68, 60, 54], which defines the task as learning user preferences and routines from historical check-ins to recommend future POIs.
C.1 Problem Formulation
Let $\mathcal{U}=\{u_{1},u_{2},\dots,u_{M}\}$ denote the set of users, $\mathcal{P}=\{p_{1},p_{2},\dots,p_{N}\}$ the set of Points of Interest (POIs), and $\mathcal{T}=\{t_{1},t_{2},\dots,t_{K}\}$ the set of timestamps, where $M,N,K\in\mathbb{N}$.
POI Definition
Each POI $p\in\mathcal{P}$ is represented as a tuple:
$p=\langle\phi,\lambda,\kappa,\alpha,\beta,\gamma\rangle,$
where:
  • $\phi$ and $\lambda$ are the latitude and longitude,
  • $\kappa$ is the POI category (e.g., restaurant, park),
  • $\alpha$ is the unique POI identifier,
  • $\beta$ is the textual address, and
  • $\gamma$ is the POI name.
Check-in Definition
A check-in is a tuple $c=\langle u,p,t\rangle\in\mathcal{U}\times\mathcal{P}\times\mathcal{T}$, indicating that user $u$ visited POI $p$ at timestamp $t$.
Trajectory Definition
A trajectory for user $u$ is defined as a temporally ordered sequence of check-ins within a fixed time interval $\delta\tau=8$ hours. Each trajectory $T^{i}_{u}(t)$ up to timestamp $t$ is defined as:
$T^{i}_{u}(t)=\{(p_{1},t_{1}),(p_{2},t_{2}),\dots,(p_{k},t_{k})\}$
such that $t_{1}<t_{2}<\dots<t_{k}=t$ and $t_{k}-t_{k-1}\leq\delta\tau$. Given a set of historical trajectories
$\mathcal{T}_{u}=\{T^{1}_{u},T^{2}_{u},\dots,T^{L}_{u}\}$
for user $u$, where $L$ is the number of such historical trajectories, the goal is to recommend the POIs that $u$ is most likely to visit next after the current contextual trajectory $T^{\prime}_{u}(t)$.
POI Recommendation Task Definition
Given a current contextual trajectory $T^{\prime}_{u}(t)$ of user $u$ up to time $t$, along with their historical trajectories $\mathcal{T}_{u}$, the task of next POI recommendation is to rank all candidate POIs $p_{i}\in\mathcal{P}$ according to the model’s predicted probability that user $u$ will visit each POI next.
Formally, the model learns a ranking function:
$f:(T^{\prime}_{u}(t),\mathcal{T}_{u})\rightarrow\{\hat{y}_{i}\}_{i=1}^{|\mathcal{P}|}$
where $\hat{y}_{i}$ denotes the predicted likelihood that user $u$ will visit POI $p_{i}$ next. Based on these scores, a ranked list of POIs is returned as recommendations.
This formulation enables POI recommendation, where the goal is to suggest a set of likely POIs that a user may visit next, based on their historical check-ins and inferred preferences. Our evaluation metrics, Acc@k and NDCG@k, assess whether the ground-truth POI appears among the top-$k$ ranked candidates, reflecting the quality of the recommended set. In particular, Acc@1 captures the stricter task of immediate next POI prediction, measuring whether the top-ranked POI matches the user’s actual next visit.
C.2 Experiment and Implementation Details
For training and evaluation, we used the LibCity [^1] https://github.com/libcity/bigscity-libcity-datasets/ library [41], which provides implementations of classical baselines including FPMC [32], RNN [41], LSTPM [35], and DeepMove [8]. The training hyperparameters are listed in Table 7 and, unless otherwise noted, follow the default configurations provided by LibCity.
For GETNext [^2] https://github.com/songyangme/GETNext [60] and STHGCN [^3] https://github.com/alipay/Spatio-Temporal-Hypergraph-Model [54], we adapted the original source code released by the respective authors. Due to variations in dataset sizes and training costs across cities, we applied different hyperparameters for some cities, as detailed in Table 8.
Table 7: Hyperparameters for Markov-based methods and recurrent networks baselines.
Hyperparameter FPMC RNN LSTPM DeepMove
Batch Size 20 20 20 20
Learning Rate 5e-4 1e-3 1e-4 1e-3
Max Epoch 1 30 40 30
Location Embedding Size 64 500 500 500
Hidden Embedding Size N/A 500 500 500
Dropout N/A 0.3 0.8 0.5
Table 8: Hyperparameters for Transformer-based graph neural networks.
Model Cities Batch Size LR Epochs
GETNext Beijing, Melbourne, Moscow, Shanghai, Sydney, Tokyo 16 1e-3 200
Istanbul, Kuwait City, New York, Petaling Jaya, São Paulo 16 1e-4 20
Jakarta 16 5e-5 20
STHGCN Beijing, Melbourne, Shanghai, Sydney, Tokyo 16 1e-4 20
Istanbul, Jakarta, Kuwait City, Moscow, New York, Petaling Jaya, São Paulo 64 1e-4 20
All modified code implementations are available as submodules in our main dataset repository. Experiments were conducted using GPUs provided by the School of Computer Science and Engineering at UNSW Sydney via the Wolfpack computational cluster, which includes NVIDIA L4, L40S, and H100 GPUs.
C.3 Supplementary Results
We report the full results of our supervised POI recommendation baselines in Table 9 and 10, using three evaluation metrics: Acc@1, Acc@5, and NDCG@5.
Table 9: Performance of supervised POI recommendation baselines across 6 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, and Moscow. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model Beijing Istanbul Jakarta Kuwait City Melbourne Moscow
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.000 0.021 0.009 0.026 0.074 0.050 0.029 0.085 0.058 0.021 0.089 0.054 0.062 0.147 0.107 0.059 0.129 0.094
RNN 0.085 0.183 0.134 0.077 0.178 0.130 0.049 0.115 0.083 0.087 0.203 0.146 0.059 0.105 0.083 0.075 0.164 0.122
LSTPM 0.127 0.211 0.169 0.142 0.286 0.217 0.099 0.210 0.157 0.180 0.362 0.275 0.091 0.204 0.150 0.151 0.300 0.229
DeepMove 0.106 0.261 0.190 0.150 0.298 0.228 0.103 0.212 0.160 0.179 0.360 0.274 0.083 0.179 0.134 0.143 0.283 0.217
GETNext 0.433 0.527 0.486 0.146 0.268 0.210 0.155 0.257 0.209 0.175 0.322 0.251 0.100 0.250 0.179 0.175 0.335 0.260
STHGCN 0.453 0.640 0.552 0.241 0.385 0.318 0.197 0.334 0.270 0.225 0.394 0.314 0.168 0.318 0.247 0.223 0.382 0.308
Table 10: Performance of supervised POI recommendation baselines across 6 cities: New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.032 0.090 0.061 0.026 0.084 0.057 0.030 0.079 0.055 0.084 0.154 0.120 0.075 0.180 0.131 0.176 0.291 0.239
RNN 0.061 0.119 0.092 0.064 0.148 0.107 0.097 0.191 0.147 0.055 0.120 0.090 0.080 0.164 0.125 0.133 0.254 0.197
LSTPM 0.099 0.206 0.155 0.099 0.222 0.163 0.158 0.319 0.243 0.099 0.195 0.149 0.141 0.265 0.206 0.225 0.394 0.315
DeepMove 0.097 0.195 0.149 0.112 0.234 0.175 0.160 0.310 0.240 0.085 0.168 0.128 0.129 0.240 0.188 0.201 0.362 0.288
GETNext 0.134 0.263 0.202 0.139 0.254 0.200 0.202 0.360 0.286 0.115 0.230 0.177 0.181 0.347 0.266 0.180 0.361 0.275
STHGCN 0.146 0.259 0.207 0.174 0.301 0.241 0.250 0.425 0.344 0.193 0.329 0.264 0.227 0.378 0.307 0.250 0.432 0.350
Appendix D Analyzing Urban Features and POI Recommendation Performance
As discussed in Section 2.2, several hypotheses have been proposed to explain why POI recommendation models perform better in certain cities than others. These hypotheses aim to uncover how various urban features affect model performance. For example, Gowalla-CA [4, 62] often yields lower accuracy compared to FSQ-NYC and FSQ-TKY [57], suggesting that some cities may be inherently harder to model. In this analysis, we focus on supervised models only.
Prior studies [60, 54, 22] have suggested several features as potential explanatory variables, including:
  • Number of unique check-ins,
  • Number of unique trajectories,
  • Number of unique POI categories,
  • Geographical area (larger areas are assumed to be harder to model), and
  • POI density or spatial sparsity (i.e., unique POIs per unit area).
We also propose several additional features for consideration:
  • Number of unique POIs,
  • Check-in density (unique check-ins per area),
  • Trajectory density (unique trajectories per area), and
  • Category entropy, our proposed feature capturing category diversity.
Category entropy, based on Shannon entropy, measures how evenly POI categories are distributed in a city. A higher entropy suggests that check-ins are spread across a wide variety of categories, while a lower entropy indicates a concentration in fewer types. The formula for Shannon entropy is:
$H=-\sum_{i=1}^{N}p_{i}\log(p_{i})$ (1)
where $p_{i}$ is the proportion of venues in category $i$, and $N$ is the total number of POI categories. The proportion $p_{i}$ is defined as:
$p_{i}=\frac{c_{i}}{\sum_{j=1}^{N}c_{j}}$ (2)
where $c_{i}$ is the count of venues in category $i$. In other words, $p_{i}$ represents the fraction of all venues that belong to category $i$.
Moreover, previous studies have primarily focused on only three datasets: FSQ-NYC, FSQ-TKY, and Gowalla-CA. In contrast, Massive-STEPS provides broader coverage across 12 cities, enabling a more comprehensive and robust analysis. To examine the relationship between urban features and model performance, we averaged the three evaluation metrics across six supervised baselines for each city and computed the Spearman correlation with each candidate feature. To further support our findings, we also included the results of GETNext [60] and STHGCN [54] on FSQ-NYC, FSQ-TKY, and Gowalla-CA, calculated their corresponding urban features, and averaged the reported metrics of each city. Fig. 6 presents the correlations between all nine features and the average performance metric.
Among all features, category entropy shows the strongest correlation with model performance, with a Spearman correlation of $r=-0.736$ ($p=0.002$). This suggests that cities with more evenly distributed POI categories tend to yield lower prediction accuracy. Intuitively, when no single category dominates (a city has roughly equal proportions of restaurants, cafes, homes, and other POIs), it becomes more difficult for models to predict a user’s next destination. In these cases, user behavior is more varied and less predictable. On the other hand, cities with more skewed category distributions (e.g., mostly food places or mostly residential areas) tend to have more consistent patterns of movement, making them easier for models to learn and predict. Interestingly, our finding contradicts the hypothesis proposed by LLM4POI [22], which suggests that FSQ-NYC is easier to model than Gowalla-CA due to the former’s vast number of POI categories, which were supposed to provide richer contextual signals for the model. Our results indicate that it is not the number of categories that matters, but rather how these categories are distributed.
Figure 6: Spearman correlation between nine candidate urban features and the mean score of POI recommendation models across 15 cities, including Massive-STEPS (ours), FSQ [57], and Gowalla [4, 62].
Appendix E Zero-shot POI Recommendation: Task Details
E.1 Problem Formulation
The zero-shot POI recommendation task follows the same problem formulation as its supervised counterpart (see Section C.1). The key difference is that in this setting, the model parameters remain frozen and the models are pre-trained, rather than trained from randomly initialized weights.
E.2 Experiment and Implementation Details
Preprocessing
We adopted the AgentMove [^4] https://github.com/tsinghua-fib-lab/agentmove/ library [7], which provides implementations of three LLM methods: LLM-Mob [42], LLM-ZS [1], and LLM-Move [10]. The preprocessing steps used by AgentMove are as follows.
First, we selected 200 random users from the test set and sampled one random trajectory for each user. This trajectory serves as the context stays, representing the current trajectory to be predicted. The historical stays are composed of the most recent 15 trajectories from the same user, drawn from the training set. Each check-in is described by four attributes: the hour (in 12-hour format), the day of the week, the POI ID, and the POI category name.
Second, the LLMs are set to return outputs in JSON format, generating the top 5 predicted POI IDs along with an explanation of their reasoning. Following the AgentMove setup and to ensure replicability, we set the generation parameters as follows: a temperature of 0.0, a maximum output length of 1000 tokens, and an input context window capped at 2000 tokens.
Prompting
Prompt templates for each method, LLM-Mob, LLM-ZS, and LLM-Move, are presented in Listing 1, 2, and 3, respectively.
1Your task is to predict a users next location based on his/her activity pattern.
2You will be provided with <history> which is a list containing this users historical stays, then <context> which provide contextual information
3about where and when this user has been to recently. Stays in both <history> and <context> are in chronological order.
4Each stay takes on such form as (start_time, day_of_week, duration, place_id). The detailed explanation of each element is as follows:
5start_time: the start time of the stay in 12h clock format.
6day_of_week: indicating the day of the week.
7duration: an integer indicating the duration (in minute) of each stay. Note that this will be None in the <target_stay> introduced later.
8place_id: an integer representing the unique place ID, which indicates where the stay is.
9
10Then you need to do next location prediction on <target_stay> which is the prediction target with unknown place ID denoted as <next_place_id> and
11unknown duration denoted as None, while temporal information is provided.
12
13Please infer what the <next_place_id> might be (please output the 10 most likely places which are ranked in descending order in terms of probability), considering the following aspects:
141. the activity pattern of this user that you learned from <history>, e.g., repeated visits to certain places during certain times;
152. the context stays in <context>, which provide more recent activities of this user;
163. the temporal information (i.e., start_time and day_of_week) of target stay, which is important because peoples activity varies during different time (e.g., nighttime versus daytime)
17and on different days (e.g., weekday versus weekend).
18
19Please organize your answer in a JSON object containing following keys:
20"prediction" (the ID of the five most probable places in descending order of probability) and "reason" (a concise explanation that supports your prediction). Do not include line breaks in your output.
21
22The data are as follows:
23<history>: {historical_stays}
24<context>: {context_stays}
25<target_stay>: {target_time, target_day_of_week}
Listing 1: Prompt for LLM-Mob
1Your task is to predict <next_place_id> in <target_stay>, a location with an unknown ID, while temporal data is available.
2
3Predict <next_place_id> by considering:
41. The users activity trends gleaned from <historical_stays> and the current activities from <context_stays>.
52. Temporal details (start_time and day_of_week) of the target stay, crucial for understanding activity variations.
6
7Present your answer in a JSON object with:
8"prediction" (IDs of the five most probable places, ranked by probability) and "reason" (a concise justification for your prediction).
9
10The data:
11<historical_stays>: {historical_stays}
12<context_stays>: {context_stays}
13<target_stay>: {target_time, target_day_of_week}
Listing 2: Prompt for LLM-ZS
1<long-term check-ins> [Format: (POIID, Category)]: {historical_stays}
2<recent check-ins> [Format: (POIID, Category)]: {context_stays}
3<candidate set> [Format: (POIID, Distance, Category)]: {candidates}
4Your task is to recommend a users next point-of-interest (POI) from <candidate set> based on his/her trajectory information.
5The trajectory information is made of a sequence of the users <long-term check-ins> and a sequence of the users <recent check-ins> in chronological order.
6Now I explain the elements in the format. "POIID" refers to the unique id of the POI, "Distance" indicates the distance (kilometers) between the user and the POI, and "Category" shows the semantic information of the POI.
7
8Requirements:
91. Consider the long-term check-ins to extract users long-term preferences since people tend to revisit their frequent visits.
102. Consider the recent check-ins to extract users current perferences.
113. Consider the "Distance" since people tend to visit nearby pois.
124. Consider which "Category" the user would go next for long-term check-ins indicates sequential transitions the user prefer.
13
14Please organize your answer in a JSON object containing following keys:
15"prediction" (10 distinct POIIDs of the ten most probable places in <candidate set> in descending order of probability), and "reason" (a concise explanation that supports your recommendation according to the requirements). Do not include line breaks in your output.
Listing 3: Prompt for LLM-Move
Models and Implementations
We use the following LLMs in our experiments:
All open-source models are quantized using AWQ [25] and served via vLLM [21]. Inference of open-source models was conducted on NVIDIA A100 GPUs provided by SaturnCloud [^8] https://saturncloud.io/ through the NVIDIA Academic Grant Program. We accessed Gemini via the official API and through the Google Cloud Research Credits program. All modified code implementations are publicly available in our main dataset repository.
E.3 Supplementary Results
We provide the full results of our zero-shot POI recommendation results in Table 11 and 12, providing three metrics: Acc@1, Acc@5, and NDCG@5.
Table 11: Performance of zero-shot POI recommendation baselines across 6 cities: Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, and Moscow. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Method Model Beijing Istanbul Jakarta Kuwait City Melbourne Moscow
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.115 0.308 0.226 0.080 0.225 0.160 0.100 0.245 0.174 0.095 0.270 0.185 0.060 0.150 0.108 0.130 0.245 0.187
Qwen 2.5 7B 0.058 0.385 0.218 0.035 0.240 0.148 0.105 0.245 0.179 0.080 0.220 0.155 0.030 0.130 0.083 0.090 0.270 0.185
Llama 3.1 8B 0.000 0.000 0.000 0.020 0.110 0.065 0.055 0.150 0.104 0.030 0.100 0.066 0.010 0.065 0.040 0.030 0.100 0.068
Gemma 2 9B 0.115 0.288 0.206 0.075 0.200 0.146 0.105 0.240 0.178 0.080 0.210 0.150 0.055 0.150 0.108 0.100 0.240 0.176
LLM-ZS Gemini 2 Flash 0.058 0.385 0.246 0.090 0.235 0.166 0.110 0.250 0.188 0.080 0.245 0.167 0.065 0.160 0.115 0.125 0.300 0.217
Qwen 2.5 7B 0.038 0.404 0.237 0.040 0.235 0.141 0.065 0.250 0.161 0.050 0.220 0.140 0.040 0.155 0.100 0.080 0.260 0.176
Llama 3.1 8B 0.077 0.346 0.221 0.040 0.225 0.137 0.045 0.200 0.126 0.060 0.210 0.137 0.040 0.155 0.101 0.080 0.270 0.183
Gemma 2 9B 0.096 0.308 0.217 0.045 0.225 0.141 0.105 0.250 0.180 0.070 0.230 0.153 0.050 0.140 0.100 0.080 0.290 0.194
LLM-Move Gemini 2 Flash 0.096 0.346 0.218 0.205 0.385 0.289 0.295 0.405 0.350 0.220 0.380 0.295 0.225 0.325 0.275 0.220 0.400 0.316
Qwen 2.5 7B 0.192 0.346 0.280 0.175 0.270 0.226 0.115 0.225 0.169 0.160 0.285 0.227 0.110 0.220 0.165 0.230 0.310 0.274
Llama 3.1 8B 0.058 0.135 0.100 0.015 0.055 0.036 0.015 0.025 0.021 0.010 0.035 0.023 0.040 0.195 0.123 0.005 0.065 0.031
Gemma 2 9B 0.096 0.365 0.229 0.100 0.200 0.155 0.235 0.290 0.266 0.120 0.275 0.202 0.115 0.275 0.199 0.110 0.245 0.185
Table 12: Performance of zero-shot POI recommendation baselines across 6 cities: New York, Petaling Jaya, São Paulo, Shanghai, Sydney, and Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Method Model New York Petaling Jaya São Paulo Shanghai Sydney Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.095 0.175 0.136 0.090 0.220 0.160 0.130 0.305 0.223 0.055 0.160 0.111 0.060 0.160 0.112 0.140 0.320 0.238
Qwen 2.5 7B 0.070 0.185 0.131 0.030 0.195 0.116 0.090 0.290 0.188 0.040 0.170 0.108 0.035 0.145 0.091 0.110 0.350 0.243
Llama 3.1 8B 0.025 0.090 0.061 0.010 0.090 0.050 0.030 0.165 0.098 0.005 0.020 0.013 0.020 0.085 0.053 0.005 0.045 0.025
Gemma 2 9B 0.070 0.175 0.124 0.055 0.185 0.122 0.085 0.230 0.162 0.050 0.150 0.104 0.030 0.130 0.086 0.145 0.345 0.255
LLM-ZS Gemini 2 Flash 0.080 0.170 0.129 0.110 0.210 0.164 0.150 0.315 0.235 0.065 0.160 0.113 0.060 0.155 0.111 0.160 0.380 0.278
Qwen 2.5 7B 0.050 0.180 0.116 0.045 0.175 0.111 0.095 0.290 0.198 0.045 0.155 0.103 0.045 0.170 0.109 0.120 0.365 0.257
Llama 3.1 8B 0.055 0.160 0.111 0.030 0.205 0.123 0.030 0.280 0.159 0.060 0.165 0.116 0.040 0.185 0.110 0.110 0.415 0.269
Gemma 2 9B 0.075 0.185 0.129 0.065 0.185 0.126 0.075 0.300 0.192 0.050 0.165 0.112 0.045 0.155 0.103 0.110 0.395 0.263
LLM-Move Gemini 2 Flash 0.235 0.415 0.325 0.210 0.335 0.273 0.285 0.415 0.350 0.170 0.270 0.221 0.230 0.420 0.331 0.250 0.470 0.368
Qwen 2.5 7B 0.120 0.255 0.188 0.135 0.175 0.155 0.155 0.235 0.199 0.095 0.165 0.133 0.125 0.280 0.205 0.250 0.360 0.312
Llama 3.1 8B 0.035 0.130 0.084 0.040 0.060 0.049 0.045 0.045 0.045 0.020 0.040 0.030 0.055 0.220 0.141 0.030 0.060 0.046
Gemma 2 9B 0.115 0.245 0.183 0.175 0.235 0.208 0.195 0.300 0.252 0.105 0.150 0.128 0.125 0.370 0.254 0.130 0.305 0.225
Appendix F License and Data Usage
Our work does not involve the collection of new data. Instead, we derive our resulting dataset by combining and aligning two publicly available datasets, both of which are distributed under permissive licenses. We did not scrape data from the internet or use proprietary APIs to construct this dataset.
We accessed the Semantic Trails Dataset [29] via Figshare: https://doi.org/10.6084/m9.figshare.7429076.v2. The dataset is licensed under the Creative Commons CC0 1.0 license (https://creativecommons.org/publicdomain/zero/1.0/), which allows unrestricted copying, modification, and redistribution for any purpose, including commercial use, without requiring permission.
We accessed the Foursquare Open Source Places dataset via Hugging Face: https://huggingface.co/datasets/foursquare/fsq-os-places. Foursquare Open Source Places is licensed under the Apache License, Version 2.0:
Copyright 2024 Foursquare Labs, Inc. All rights reserved. Licensed under the Apache License, Version 2.0 (the "License"); you may not use this file except in compliance with the License. You may obtain a copy of the License at:  http://www.apache.org/licenses/LICENSE-2.0 Unless required by applicable law or agreed to in writing, software distributed under the License is distributed on an "AS IS" BASIS, WITHOUT WARRANTIES OR CONDITIONS OF ANY KIND, either express or implied. See the License for the specific language governing permissions and limitations under the License.
More details are available in Foursquare’s documentation: https://docs.foursquare.com/data-products/docs/access-fsq-os-places.
We released our Massive-STEPS dataset under the same Apache Version 2.0 License, and have included Foursquare Open Source Places’ license in our hosted dataset’s README file.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims and contributions in the abstract and introduction have been clearly stated and accurately match the experimental results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discussed the limitations of the work in the dedicated “Limitations” section (see Section 5), which discussed the limitations of our dataset and benchmarks.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We ensured reproducibility by (1) releasing the dataset and all accompanying code to replicate our supervised and zero-shot POI recommendation experiments, and (2) providing detailed descriptions of preprocessing procedures, model configurations, training setups, and evaluation protocols throughout the paper (see Section C.2 and Section E.2).
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The paper provides open access to the Massive-STEPS dataset and all source code, along with detailed instructions for data access, preprocessing, and experiment reproduction. This includes scripts for preparing raw and resultant aligned data, reproducing all supervised and zero-shot POI recommendation benchmarks, and evaluating both proposed methods and baselines.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specifies all relevant training and testing details (see Section C and E). For supervised models, we describe how hyperparameters were selected (following original papers), and for zero-shot LLM baselines, we detail the prompt formats, generation parameters, and input constraints.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We did not conduct repeated experiments to calculate error bars in our benchmarking due to computational resource constraints. However, we included the p-value of our proposed hypothesis in Section D.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We stated the computational resources used and needed to reproduce our experiments in Section C.2 and Section E.2.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have reviewed and ensured that our research conforms to the NeurIPS Code of Ethics. The original datasets we used and on which we based our dataset have been anonymized to remove PII and have a clear license (see Section F).
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have discussed the positive societal benefits of POI recommendation research in Section 1 and the issues of the field, such as the existence of geographical bias in LLMs (see Section 2.2), which may cause fairness issues.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: While we did not scrape a dataset from the internet and instead, based our dataset on two existing datasets, the latter have been anonymized to remove PII and do not impose a high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the original paper that produced the Semantic Trails dataset and have clearly mentioned the license of Semantic Trails (CC0 1.0 Universal) and Foursquare Open Source Places (Apache 2.0), providing the URL. See Section F.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We released the dataset and all accompanying code with documentation to replicate our experiments.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components, and was only used for writing, editing, or formatting purposes.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.