### v3

# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins

Abstract.
Understanding human mobility through Point-of-Interest (POI) trajectory modeling is increasingly important for applications such as urban planning, personalized services, and generative agent simulation. However, progress in this field is hindered by two key challenges: the over-reliance on older datasets from 2012-2013 and the lack of reproducible, city-level check-in datasets that reflect diverse global regions. To address these gaps, we present Massive-STEPS (Massive Semantic Trajectories for Understanding POI Check-ins), a large-scale, publicly available benchmark dataset built upon the Semantic Trails dataset and enriched with semantic POI metadata. Massive-STEPS spans 15 geographically and culturally diverse cities and features more recent (2017-2018) and longer-duration (24 months) check-in data than prior datasets. We benchmarked a wide range of POI models on Massive-STEPS using both supervised and zero-shot approaches, and evaluated their performance across multiple urban contexts. By releasing Massive-STEPS, we aim to facilitate reproducible and equitable research in human mobility and POI trajectory modeling. The dataset and benchmarking code are available at: https://github.com/cruiseresearchgroup/Massive-STEPS.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
Importance of Human Mobility Data and Modeling
Human mobility data and POI trajectory modeling are essential for understanding how individuals interact with and move through physical spaces. This understanding enables a wide range of applications, including urban planning (Yuan et al., 2025), travel service recommendations (Feng et al., 2025), improved commercial advertising strategies (Yang et al., 2022b), and Point-of-Interest (POI) recommendation (Ding et al., 2020; Li et al., 2024; Zhang et al., 2025). Recently, human mobility data has become even more crucial with the increasing use of large language model (LLM) agents to simulate human-like behavior and routines (Zhou et al., 2024; Jiawei et al., 2024). However, while simulated and aggregated human mobility data are starting to gain popularity (Feng et al., 2020; Qin et al., 2023; Stanford et al., 2024; Jiang et al., 2025), they may not accurately reflect real-world, fine-grained individual human behavior (Salim et al., 2020), highlighting the value of evaluating on real-world data. These advancements are enabled by and large with Location-based Social Networks (LBSNs), which generate vast amounts of spatio-temporal data through user check-ins (Zhang et al., 2025; Li et al., 2024). This rich data source has allowed the development of POI recommendation systems that leverage users’ historical visiting behaviors to suggest relevant locations. Such systems enhance user engagement through personalization and provide commercial value to both users and businesses by aligning recommendations with individual preferences and available services (Ding et al., 2020).
Literature Gaps
Our paper addresses three critical gaps in POI trajectory modeling research and datasets. First, as shown in Fig. 7, the field remains dominated by studies focused on just two cities, New York and Tokyo, based on the Foursquare dataset curated by Yang et al. (2014). This dataset, collected in 2012-2013, raises concerns about its temporal quality, as many POIs may no longer exist and user behavior may have changed (Yeow et al., 2021). While some recent studies have expanded to other cities (Zhang et al., 2024a; Merinov and Ricci, 2024; Feng et al., 2025), they often rely on the Global-scale Check-in Dataset (GSCD) (Yang et al., 2015, 2016), which, despite its extensive coverage, is also from 2012-2013 and contains nearly 50% erroneous entries (Monti et al., 2018). Second, most existing studies are difficult to reproduce, either due to the lack of clearly defined geographic boundaries or the unavailability of the datasets themselves, hindering fair comparison and replication. Finally, we join recent efforts (Yuan et al., 2025) in advocating for the inclusion of low-resource and underrepresented cities. Expanding beyond well-studied urban centers is essential for building more generalizable and universally applicable POI models. Table 1 summarizes these limitations in terms of geographic coverage, temporal span, and reproducibility.
Figure 1. Massive-STEPS Dataset and Benchmark Overview.
Massive-STEPS Dataset and Benchmark Overview.
Massive-STEPS Dataset
In this paper, we introduced the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) Dataset, derived from the Semantic Trails dataset (STD) (Monti et al., 2018). Massive-STEPS includes high-quality check-ins from 2012-2013 and 2017-2018, providing more modern and updated POI check-in data. This supports longitudinal POI trajectory modeling studies and addresses the limitations of older datasets commonly used in prior studies. The dataset covers 15 diverse cities across multiple regions, including East, West, and Southeast Asia, North and South America, Australia, the Middle East, and Europe. Notably, we placed a deliberate emphasis on under-explored regions by including cities such as Jakarta, Kuwait City, and Petaling Jaya, filling a key gap in POI trajectory research that has largely focused on major urban centers. We further enriched STD by aligning it with Foursquare’s Open Source Places dataset, incorporating missing metadata such as geographic coordinates, POI names, and addresses.
Benchmark Tasks
To demonstrate the utility of this dataset, we conducted an extensive benchmark on three tasks as shown in Fig. 1: (1) supervised POI recommendation, (2) zero-shot POI recommendation, and (3) spatiotemporal classification and reasoning. Our benchmark covers a wide range of models, including traditional approaches, deep learning-based models, and more recent LLM-based methods. The goal of POI recommendation task is to predict a set of POIs that a user is likely to visit based on their current check-in trajectory and historical behavior. This reflects real-world applications such as personalized POI recommendations in location-based services. Similarly, the goal of spatiotemporal classification and reasoning is to assess how effectively models (e.g., LLMs) leverage, interpret, and reason about POI trajectories. In addition, the scale of our dataset allows us to examine how urban features influence POI modeling accuracy. Building on prior hypotheses, we propose a new insight: cities with more evenly distributed POI categories tend to be harder to model, as the absence of a dominant POI category makes user behavior less predictable.
Contribution
This paper introduces the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset, addressing gaps in existing POI trajectory modeling research. Current POI check-in datasets are often only from 2012-2013, skewed to a few cities, and lack semantic metadata, hindering the development of robust and globally applicable models. While datasets like GSCD and STD offer broad geographic coverage, they either suffer from an older timespan, contain erroneous data, or have missing information. Massive-STEPS overcomes these issues by providing high-quality check-ins from 2012-2013 and 2017-2018, improving temporal quality for longitudinal POI trajectory modeling studies. The dataset spans 15 diverse cities across multiple regions, with a focus on low-resource cities overlooked in previous research. Additionally, Massive-STEPS is enriched with metadata through alignment with Foursquare’s Open Source Places, providing crucial details such as POI geographical coordinates, POI names, and addresses. We also conducted an extensive benchmark on both supervised and zero-shot POI recommendation and trajectory classification tasks, evaluating a wide range of models, including traditional methods, deep learning approaches, and recent LLM-based techniques. We further analyzed which urban features affect POI modeling accuracy and found that cities with no dominant POI category tend to be harder to predict. By releasing this dataset and benchmark code publicly, we facilitate open and reproducible research, enabling future advancements in urban mobility studies.
Table 1. Comparison of check-in datasets commonly used for POI modeling tasks.
Dataset Scale xxCompletenessxx Usability
#cities Years #months POI Attributes Replicable Open
NYC/TKY (Yang et al., 2014) 2 2012-2013 11   Coordinates, Category
Gowalla-CA (Yuan et al., 2013) 1 2009-2010 21   Coordinates, Category
AgentMove (Feng et al., 2025) 12 2012-2013 17   Coordinates, Category
Massive-STEPS 15 2012-2013, 2017-2018 24 Coordinates, Category, Name, Address
2. Existing Datasets
A survey conducted by (Zhang et al., 2025) outlines the landscape of POI trajectory modeling research, covering a wide range of models and architectures used in prior studies. While it offers a high-level overview of the datasets used, it lacks a dedicated discussion or evaluation of POI datasets. We address this gap by analyzing commonly used datasets and positioning our dataset within this context.
Table 2. Overview of POI dynamics: total POIs ever opened (according to Foursquare OS Places), POIs confirmed closed by 2025, and POIs closed during 2014-2016, corresponding to the temporal gap in our dataset.
City POIs Opened Total POIs Closed Closed within 2014-2016
New York 49,218 13,009 (26.43%) 3,118 (6.34%)
Melbourne 7,699 1,850 (24.03%) 209 (2.71%)
Sydney 8,986 1,759 (19.57%) 253 (2.82%)
Moscow 17,822 3,021 (16.95%) 868 (4.87%)
São Paulo 38,377 4,990 (13.00%) 1,257 (3.28%)
Shanghai 4,462 661 (14.81%) 81 (1.82%)
Tokyo 4,725 421 (8.91%) 0 (0.00%)
Petaling Jaya 60,158 4,186 (6.96%) 1,533 (2.55%)
Istanbul 53,812 2,833 (5.26%) 481 (0.89%)
Beijing 1,127 56 (4.97%) 10 (0.89%)
Jakarta 76,116 3,527 (4.63%) 483 (0.63%)
Bandung 29,026 1,053 (3.63%) 182 (0.63%)
Palembang 4,343 143 (3.29%) 23 (0.53%)
Tangerang 12,956 383 (2.96%) 50 (0.39%)
Kuwait City 17,180 161 (0.94%) 22 (0.13%)
LBSN Check-in Data Sources
Building on the table from (Zhang et al., 2025), we investigated which datasets are most commonly used in prior studies. From their original table (Table IV), we filtered entries pertaining specifically to POI and next POI recommendation tasks and identified (1) the most frequently used LBSN check-in data sources and (2) the most commonly studied cities. As shown in Fig. 7, Foursquare remains the dominant source of LBSN data in existing studies, appearing in almost 50% of the surveyed works. While several variants of Foursquare datasets have been employed, the most widely used are the NYC and Tokyo Dataset (Yang et al., 2014) (often abbreviated as FSQ-NYC and FSQ-TKY) and the Global-scale Check-in Dataset (GSCD) (Yang et al., 2015, 2016), curated by the same authors. Other LBSN sources occasionally used include Gowalla (Cho et al., 2011), Brightkite (Cho et al., 2011), and Weeplaces (Liu et al., 2017).
Saturated to Two Cities and Old Timespan
Due to the widespread use of FSQ-NYC and FSQ-TKY (Yang et al., 2014), the majority of POI trajectory studies are disproportionately focused on these two cities, as illustrated in Fig. 7. While there is nothing inherently problematic about NYC and Tokyo, there has been growing interest in expanding research to a broader range of cities, particularly those that are underexplored or considered low-resource (Yuan et al., 2025), as cultural and regional differences influence collective mobility behaviors. For instance, in some cities, residents tend to commute to business districts in the morning, whereas in others, nightlife activities such as visiting bars after work are more common (Yang et al., 2015). Ensuring diverse geographic coverage is increasingly important, especially as LLMs are adopted for POI trajectory modeling tasks. LLMs are known to exhibit geographical biases against regions with lower socioeconomic conditions (Manvi et al., 2024). Whether LLMs can generalize across diverse urban environments is to be investigated.
In addition, because many studies rely on the FSQ-NYC and FSQ-TKY, they are often constrained to the timespan it covers: check-in data from 2012 to 2013. However, POI data is inherently dynamic: venues may have closed, relocated, or changed in category over time. Table 2 illustrates this by detailing the total number of POIs ever opened, the number of POIs closed by 2025, and the number of POIs that closed between 2014 and 2016 alone. These substantial closure rates in major cities highlight the critical need for newer, multi-period datasets that reflect the actual lifecycle of POIs.
Yeow et al. (2021) further underscores the importance of validating the temporal quality of POI datasets by recording whether and when a venue’s information has been updated to reflect real-world changes. This is particularly critical, as recommender systems should avoid suggesting POIs that no longer exist or have undergone substantial changes (e.g., a former bookstore converted into a coworking space) and behave dynamically over longitudinal periods (Yabe et al., 2024). Moreover, visitation patterns may change over time depending on the city (see Fig.2), potentially due to shifts in routines. For instance, the opening of a new train station can noticeably alter commuting patterns and the popularity of nearby POIs.
Figure 2. Top 10 most visited POI categories in Bandung and Kuwait City across two time periods. Remaining cities are shown in Appendix B.
Top 10 most visited POI categories in Bandung and Kuwait City across two time periods.
Low Data Quality: Erroneous Entries
More recently, researchers have begun leveraging the broader Global-scale Check-in Dataset (GSCD) (Yang et al., 2015, 2016), which spans 415 cities across 77 countries. Despite its wider geographic coverage, GSCD is temporally limited to the same 2012-2013 period as FSQ-NYC and FSQ-TKY, and thus suffers from similar issues of temporal quality. More critically, Monti et al. (2018) demonstrated that GSCD suffers from significant data quality issues, with over 14 million check-ins (about 44%) of the dataset flagged as erroneous due to anomalous user behavior. These include (1) repeated check-ins at the same venue, (2) check-ins occurring within implausibly short time intervals (less than one minute), and (3) transitions between venues that would require travel speeds exceeding Mach 1, which are physically impossible.
To address these limitations, Monti et al. (2018) introduced the Semantic Trails Dataset (STD), which applies systematic filtering procedures to enhance data quality. STD comprises two subsets: a cleaned version of GSCD covering 2012-2013 (STD 2013), and a newer collection of check-ins from 2017-2018 (STD 2018), sourced from Foursquare Swarm. STD 2018 also spans a wider range of cities, making it valuable for capturing globally distributed user behavior, in contrast to GSCD’s focus on densely populated urban centers. Both subsets follow the same rigorous filtering criteria, resulting in a higher-quality check-in dataset for downstream POI trajectory modeling tasks. Given these improvements, we adopted STD as the source for our check-in dataset.
Poor Reproducibility
Another persistent challenge in POI trajectory research is the lack of reproducibility in dataset preprocessing. While some recent studies utilize datasets like GSCD to cover a wide range of cities, they often omit important details needed for replicating their data filtering processes. For example, (Feng et al., 2025) and (Zuo and Zhang, 2024) conducted city-level filtering, but they did not specify how the city boundaries were defined or what distance-based thresholds were used. Similarly, the Weeplaces dataset used in (Chen and Zhu, 2025) and (Cao et al., 2023) is no longer available. To further support this claim, we provide an extensive list of dataset reproducibility issues in all the studies reviewed in (Zhang et al., 2025), in Table 8. As shown, almost none of the datasets used in these works are fully reproducible or publicly available, except for FSQ-NYC/TKY (Yang et al., 2014) and Gowalla-CA (Yuan et al., 2013), leading to a heavy reliance on these datasets.
3. Massive-STEPS Dataset
3.1. Creation Process
Massive-STEPS is derived from Semantic Trails [^1] Semantic Trails is originally released with a CC0 1.0 Universal license. (Monti et al., 2018), incorporating check-ins from both the 2013 and 2018 subsets. We utilize two additional components from STD: (1) the cities metadata file, which provides the latitude and longitude of administrative regions (e.g., towns, suburbs) along with their corresponding country codes obtained from GeoNames; and (2) the POI category mapping, which links each Foursquare Category ID to its descriptive name (e.g., ”Restaurant”). Based on this metadata, each POI is thus associated with several attributes: Foursquare Place ID, Foursquare Category ID, category name, latitude/longitude of the administrative region, the administrative region name, and the country code. For anonymization and privacy purposes and model training, we apply ordinal encoding to the Place IDs and Category IDs, assigning each a unique integer index.
Trajectory Grouping
Most POI trajectory models operate on sequences of check-ins, commonly referred to as trajectories. The model is tasked with predicting the next POIs a user is likely to visit, given the current trajectory. STD conveniently provides pre-grouped trajectories (trails) by applying a time interval-based grouping: for each user, check-ins that occur within a time interval of $\delta_{\tau}=8$ hours are grouped into the same trajectory.
Matching Trajectories to Target Cities
To obtain city-specific datasets, we matched trajectories to the target cities. For each city, we obtain geographic boundaries from OpenStreetMap and retrieve its GeoJSON file via the Overpass API. The GeoJSON file contains a polygon defining the city’s boundary in latitude and longitude. Using this boundary, we filter check-ins by comparing the latitude/longitude of each POI’s administrative region and retain only those that are within the city’s polygon. This ensures that all retained trajectories are spatially grounded within the designated city. For reproducibility purposes, we also include the GeoJSON files with our dataset’s release.
Filtering Short Trajectories and Inactive Users
To ensure data quality, we apply an additional filtering step by removing trajectories with fewer than two check-ins and excluding users with fewer than three trajectories. This prevents the model from learning from overly sparse or irrelevant data.
Train, Validation, and Test Splits
We split trajectories into training, validation, and test sets in a ratio of 7:1:2, following (Feng et al., 2025). We ensure that all users in the test set appear at least once in the training or validation set, following prior studies (Yang et al., 2022c; Yan et al., 2023).
Table 3. Statistics of 15 Massive-STEPS subsets, including the number of users, trajectories, POIs, check-ins, and train, validation, and test trajectory counts. $\mu_{\text{len}}$ denotes the mean number of check-ins per trajectory, and $\mu_{\text{int}}$ denotes the mean time interval between check-ins (in hours). For comparison, we also include statistics from existing Foursquare- and Gowalla-based datasets. We report the version used in (Yan et al., 2023).
City Users Trajectories POIs Check-ins #train #val #test $\mu_{\text{len}}$ $\mu_{\text{int}}$
NYC and Tokyo Check-in Dataset (Yang et al., 2014)
New York 1,048 14,130 4,981 103,941 72,206 1,400 1,347 7.55 7.27
Tokyo 2,282 65,499 7,833 405,000 274,597 6,868 7,038 6.32 5.47
Gowalla (Cho et al., 2011; Yuan et al., 2013)
California 3,957 45,123 9,690 238,369 154,253 3,529 2,780 5.24 8.37
Massive-STEPS
Bandung 3,377 55,333 29,026 161,284 38,732 5,534 11,067 2.91 3.17
Beijing 56 573 1,127 1,470 400 58 115 2.57 3.10
Istanbul 23,700 216,411 53,812 544,471 151,487 21,641 43,283 2.52 4.36
Jakarta 8,336 137,396 76,116 412,100 96,176 13,740 27,480 3.00 2.81
Kuwait City 9,628 91,658 17,180 232,706 64,160 9,166 18,332 2.54 5.31
Melbourne 646 7,864 7,699 22,050 5,504 787 1,573 2.80 3.27
Moscow 3,993 39,485 17,822 105,620 27,639 3,949 7,897 2.67 3.36
New York 6,929 92,041 49,218 272,368 64,428 9,204 18,409 2.96 3.16
Palembang 267 4,699 4,343 14,467 3,289 470 940 3.08 3.17
Petaling Jaya 14,308 180,410 60,158 506,430 126,287 18,041 36,082 2.81 2.96
São Paulo 5,822 89,689 38,377 256,824 62,782 8,969 17,938 2.86 3.54
Shanghai 296 3,636 4,462 10,491 2,544 364 728 2.89 3.02
Sydney 740 10,148 8,986 29,900 7,103 1,015 2,030 2.95 3.33
Tangerang 1,437 15,984 12,956 45,521 11,188 1,599 3,197 2.85 3.24
Tokyo 764 5,482 4,725 13,839 3,836 549 1,097 2.52 5.16
POI Enrichment via Foursquare OS Places
Since the POIs in STD include their corresponding Foursquare Place IDs, we matched them directly with entries in the Foursquare OS Places dataset using these IDs as the key. This one-to-one ID correspondence allows for a straightforward join operation, enriching each POI with additional metadata such as its precise latitude and longitude, name (e.g., of a restaurant or subway station), and address [^2] Foursquare OS Places is originally released with an Apache 2.0 license.. However, not all POIs in the Foursquare OS Places dataset include the full metadata, particularly those categorized as private residences, which are expectedly excluded due to privacy restrictions.
3.2. Description and Addressing Literature Gaps
Massive-STEPS is a city-level POI check-in dataset consisting of real-world, user check-in trajectories from 15 cities: Bandung, Beijing, Istanbul, Jakarta, Kuwait City, Melbourne, Moscow, New York, Palembang, Petaling Jaya, São Paulo, Shanghai, Sydney, Tangerang, and Tokyo. It features anonymized POI check-ins enriched with geographical metadata to support spatiotemporal and sequential modeling tasks. City-level statistics, along with comparisons to existing datasets, are presented in Table 3. Table 9 shows the available fields in the dataset and provides an example for each field. Fig. 8 shows a world map highlighting the cities included in the dataset.
Massive-STEPS offers a more comprehensive representation of urban mobility compared to typical POI check-in datasets. As shown in Table 3, datasets like FSQ-NYC and FSQ-TKY (Yang et al., 2014) contain fewer than 10,000 candidate POI locations. In contrast, cities in Massive-STEPS cover significantly more POIs: Massive-STEPS New York has over 49,000 POIs, while Massive-STEPS Jakarta exceeds 76,000. Massive-STEPS Istanbul, one of the largest subsets, features a large user base of 23,700, offering a broad range of user behaviors. Although some Massive-STEPS subsets are smaller than their FSQ counterparts (e.g., Tokyo), we attribute this to the strict filtering procedures applied by STD to remove erroneous entries, as explained in Section 2. Analysis of the dataset reveals a long-tailed distribution in both trajectory length (Fig. 3) and user activity (Fig. 4). Most users exhibit cold-start behavior, contributing only a few short trajectories per outing. This prevalence of short and sparse trajectories underscores the necessity for models capable of learning and generalizing effectively from limited sequential data.
Massive-STEPS provides temporal coverage across two non-consecutive periods, spanning 2012-2013 and 2017-2018 (24 months in total), which enables longitudinal analysis of urban mobility and POI visitation patterns. This temporal span allows us to evaluate how POI recommendation models perform across different time periods (see Section 4.2.3). POIs are inherently dynamic, with substantial closure rates observed in major cities such as New York, Melbourne, and Sydney (see Table 2), indicating that the urban POI landscape is far from static and underscoring the importance of multi-period datasets like ours. Furthermore, we observe clear distributional shifts in the most visited POI categories between the two periods, as illustrated in Fig. 2, indicating that visitation behaviors can change substantially even within the same city
Beyond its scale, Massive-STEPS addresses the oversaturation of FSQ-NYC and FSQ-TKY in POI trajectory modeling research. Notably, Massive-STEPS includes low-resource and previously underexplored cities in human mobility studies, such as Petaling Jaya and Kuwait City, both of which are among the cities with the highest number of check-ins from STD. This broader coverage opens new research opportunities for studying location-based behaviors across diverse cultural and geographic contexts. Furthermore, since Massive-STEPS is based on STD, it benefits from the carefully filtered, high-quality check-ins and a longer, more recent timespan. These characteristics make Massive-STEPS a more relevant and reliable resource for modeling human mobility patterns.
[…]
Model Bandung Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Palembang Petaling Jaya São Paulo Shanghai Sydney Tangerang Tokyo
A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5
FPMC 0.048 0.083 0.000 0.009 0.026 0.050 0.029 0.058 0.021 0.054 0.062 0.107 0.059 0.094 0.032 0.061 0.102 0.136 0.026 0.057 0.030 0.055 0.084 0.120 0.075 0.131 0.104 0.166 0.176 0.239
RNN 0.062 0.099 0.085 0.134 0.077 0.130 0.049 0.083 0.087 0.146 0.059 0.083 0.075 0.122 0.061 0.092 0.049 0.085 0.064 0.107 0.097 0.147 0.055 0.090 0.080 0.125 0.087 0.135 0.133 0.197
[…]
Flashback 0.103 0.155 0.278 0.359 0.137 0.200 0.113 0.166 0.150 0.227 0.079 0.115 0.121 0.180 0.086 0.123 0.135 0.198 0.072 0.119 0.159 0.222 0.096 0.132 0.127 0.174 0.185 0.262 0.170 0.237
Graph-Flashback 0.091 0.146 0.304 0.373 0.137 0.202 0.072 0.121 0.115 0.181 0.093 0.143 0.125 0.188 0.059 0.093 0.148 0.208 0.072 0.119 0.096 0.156 0.111 0.151 0.135 0.188 0.200 0.292 0.180 0.259
LoTNext 0.076 0.125 0.243 0.320 0.105 0.168 0.044 0.078 0.132 0.209 0.085 0.133 0.089 0.145 0.061 0.099 0.140 0.208 0.055 0.097 0.105 0.166 0.093 0.139 0.123 0.176 0.126 0.190 0.191 0.258
[…]
STHGCN 0.219 0.302 0.453 0.552 0.241 0.318 0.197 0.270 0.225 0.314 0.168 0.247 0.223 0.308 0.146 0.207 0.246 0.341 0.174 0.241 0.250 0.344 0.193 0.264 0.227 0.307 0.293 0.400 0.250 0.350
UniMove 0.007 0.033 0.036 0.128 0.015 0.038 0.004 0.020 0.023 0.073 0.008 0.037 0.009 0.030 0.004 0.016 0.009 0.035 0.008 0.034 0.002 0.009 0.000 0.029 0.015 0.059 0.001 0.029 0.032 0.072
Massive-STEPS is designed to be easily extended to other geographical regions. Since the data processing code is open-source and fully reproducible, adding a new city only requires its geographic boundaries from OpenStreetMap. Moreover, Massive-STEPS is scalable to higher levels of geographic granularity, enabling the creation of provincial, state, and country-level POI check-in datasets, which support collective mobility studies at broader geographic scales. To ensure full transparency and facilitate further research, all benchmarking scripts and evaluation code used in our experiments are also publicly released.
Figure 3. Distribution of trail lengths, showing a long-tailed pattern with most trajectories consisting of a few check-ins.
Distribution of trail lengths, showing a long-tailed pattern with most trajectories consisting of a few check-ins.
Figure 4. Distribution of user activity based on the number of trajectories per user, indicating a cold-start-heavy dataset.
Distribution of user activity based on the number of trajectories per user, indicating a cold-start-heavy dataset.
4. Benchmark Tasks
4.1. POI Recommendation
This benchmark focuses on POI recommendation, where the goal is to predict a user’s next visit based on their previous check-ins. The input is a trajectory of visited places, and the model is expected to suggest a set of $K$ POIs the user might visit next. It is a supervised task, trained on all available historical trajectories to learn personalized movement patterns. Appendix C provides details on problem formulation, hyperparameters, experimental setups, and full evaluation results.
4.1.1. Experimental Setup
We adopted the predefined trajectories from the original STD, where check-ins are grouped into sequences based on fixed time intervals (see Section 3.1). All input features are numerically encoded, enabling straightforward use across experiments. Models typically use four feature types: (1) social: user ID; (2) spatial: POI ID and geographic coordinates; (3) temporal: check-in timestamp; and (4) categorical: POI category. For POIs lacking exact geographic coordinates due to privacy (see Section 3.1), we defaulted to the coordinates of their administrative region as a spatial proxy. We evaluated four kinds of architectures: (1) Markov-based methods: FPMC (Rendle et al., 2010), (2) classical deep learning models: RNN (Wang et al., 2021a), LSTPM (Sun et al., 2020), DeepMove (Feng et al., 2018), Flashback (Yang et al., 2021), Graph-Flashback (Rao et al., 2022), LoTNext (Xu et al., 2024), (3) Transformer-based graph neural networks: GETNext (Yang et al., 2022c), STHGCN, and (4) trajectory foundation model: UniMove (Han et al., 2025b). We employed two commonly used metrics in POI recommender systems: Acc@k, which checks if the true POI appears in the top-k predicted results, and NDCG@k, which measures the ranking quality of the suggested results.
[…]
Method Bandung Beijing Istanbul Jakarta Kuwait City Melbourne Moscow New York Palembang Petaling Jaya São Paulo Shanghai Sydney Tangerang Tokyo
A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5 A@1 N@5
[…]
Qwen 2.5 7B 0.060 0.111 0.058 0.218 0.035 0.148 0.105 0.179 0.080 0.155 0.030 0.083 0.090 0.185 0.070 0.131 0.075 0.143 0.030 0.116 0.090 0.188 0.040 0.108 0.035 0.091 0.095 0.196 0.110 0.243
Llama 3.1 8B 0.010 0.055 0.000 0.000 0.020 0.065 0.055 0.104 0.030 0.066 0.010 0.040 0.030 0.068 0.025 0.061 0.005 0.025 0.010 0.050 0.030 0.098 0.005 0.013 0.020 0.053 0.020 0.073 0.005 0.025
[…]
Qwen 2.5 7B 0.055 0.126 0.038 0.237 0.040 0.141 0.065 0.161 0.050 0.140 0.040 0.100 0.080 0.176 0.050 0.116 0.050 0.135 0.045 0.111 0.095 0.198 0.045 0.103 0.045 0.109 0.100 0.215 0.120 0.257
Llama 3.1 8B 0.045 0.131 0.077 0.221 0.040 0.137 0.045 0.126 0.060 0.137 0.040 0.101 0.080 0.183 0.055 0.111 0.070 0.154 0.030 0.123 0.030 0.159 0.060 0.116 0.040 0.110 0.080 0.173 0.110 0.269
Gemma 2 9B 0.065 0.130 0.096 0.217 0.045 0.141 0.105 0.180 0.070 0.153 0.050 0.100 0.080 0.194 0.075 0.129 0.060 0.150 0.065 0.126 0.075 0.192 0.050 0.112 0.045 0.103 0.100 0.227 0.110 0.263
[…]
Llama 3.1 8B 0.030 0.033 0.058 0.100 0.015 0.036 0.015 0.021 0.010 0.023 0.040 0.123 0.005 0.031 0.035 0.084 0.010 0.013 0.040 0.049 0.045 0.045 0.020 0.030 0.055 0.141 0.000 0.003 0.030 0.046
[…]
Figure 5. Spearman correlation between category entropy and mean score of POI recommendation models. Comparison against other city features are provided in Appendix D.
Spearman correlation between category entropy and mean score of POI recommendation models.
4.1.2. Main Results
As shown in Table 4, STHGCN achieves the highest average Acc@1 and NDCG@5 across all cities, followed closely by GETNext, demonstrating the effectiveness of GNNs compared to classical recurrent models. The top model attained a mean Acc@1 of 23.4%, comparable to previous studies on similarly sized datasets (Feng et al., 2025). Interestingly, newer variants of Flashback, such as Graph-Flashback and LoTNext, do not always provide a significant improvement, contrary to their original claims. We note that in their respective studies, these methods were evaluated only on two datasets (Gowalla and Foursquare); our results further highlight the importance of benchmarking breadth to ensure that improvements generalize across datasets and diverse urban environments. Similarly, pre-training UniMove from scratch [^3] UniMove pre-trained weights are not publicly available. struggled to surpass older baselines. We attribute this to the high number of cold-start trajectories (see Fig. 3), as the method fails to generalize to the extremely short input sequences typical of real-world data. In these cases, next-token prediction loss struggles to provide meaningful training signal.
4.1.3. Correlating Urban Features with Model Scores
We also examined the impact of urban features on POI recommendation accuracy by computing Spearman correlations between city features and model performance. As shown in Fig. 11, we found that category entropy, based on Shannon entropy, shows a strong negative correlation with predictive accuracy (r=0.646,p<0.05r=-0.646,p<0.05). Our finding suggests that cities with more evenly distributed POI categories tend to be harder to predict. Conversely, cities dominated by a few specific POI categories, such as those with a significant density of Cafés, Restaurants, etc., tend to be easier to model and predict. Intuitively, these specialized urban environments often facilitate more regular and predictable travel behaviors, as users frequently visit these common venue types as part of their daily routines. Further details regarding the relationship between category entropy and model accuracies are provided in Appendix D.
4.2. Zero-shot POI Recommendation
This benchmark focuses on zero-shot POI recommendation via LLMs, where the goal is to predict a user’s next visit based on their previous check-ins (similar to its supervised counterpart) without additional model fine-tuning. The input is a user trajectory transformed into a textual prompt, and the model ranks a set of $K$ candidate POIs to identify the next likely destination. Appendix E provides details on problem formulation, prompts, experimental setups, and full evaluation results.
Figure 6. Acc@1 of supervised (solid lines) and LLM-Move (dashed lines) models across 15 cities.
Acc@1 of supervised (solid lines) and LLM-Move (dashed lines) models across 15 cities.
4.2.1. Experimental Setup
For zero-shot recommendation, trajectories are converted into textual prompts (Xue et al., 2022; Xue and Salim, 2024). We adapted the prompt templates from (Feng et al., 2025), which implemented the three LLM methods evaluated in this study: LLM-Mob (Wang et al., 2023c), LLM-ZS (Beneduce et al., 2024), and LLM-Move (Feng et al., 2024a). Since LLMs can leverage contextual information, features do not need numerical encoding; we used each check-in’s timestamp, POI category name, and POI ID. For a robust evaluation, we tested each method on four LLMs: one closed-source API (Gemini 2.0 Flash (Team and et al., 2024a)) and three open-source instruction-tuned models: Qwen 2.5 7B (Team, 2024), Llama 3.1 8B (Grattafiori et al., 2024), and Gemma 2 9B (Team and et al., 2024b). We used the same metrics as in the supervised setting.
4.2.2. Main Results
As shown in Table 5, LLM-Move (Feng et al., 2024a) outperformed the other two methods due to its prompt design which provides candidate POIs rather than relying solely on historical or contextual trajectories, unlike LLM-Mob and LLM-ZS. Importantly, LLM-Move selects the top $N$ closest locations from a user’s last check-in, using a distance-based heuristic echoing TimeGeo (Jiang et al., 2016). Across LLMs, Gemini 2.0 Flash achieved the highest accuracy across all prompting strategies, with Qwen 2.5 7B and Gemma 2 9B as strong open-source alternatives. Notably, as shown in Fig. 6, these zero-shot methods matched or exceeded supervised baselines in several cities (e.g., Jakarta, Kuwait City, Moscow), demonstrating their effectiveness without fine-tuning.
[…]
Time Period Model Bandung Beijing Istanbul Jakarta KC Melbourne Moscow NY Palembang PJ SP Shanghai Sydney Tangerang Tokyo
[…]
Llama 3.1 8B 0.031 0.041 0.007 0.010 0.000 0.039 0.005 0.032 0.010 0.014 0.048 0.006 0.064 0.000 N/A
[…]
4.2.3. Longitudinal Experiments and Results
To examine temporal changes in mobility patterns, we split the test sets into two periods (2012-2013 and 2017-2018) and evaluated the same four LLMs with LLM-Move (Feng et al., 2024a), which was the strongest-performing approach in our zero-shot experiments. As shown in Table 6, zero-shot accuracy generally declined in the 2017-2018 period, except for Jakarta and Tangerang, indicating that user trajectories in later years tend to be more challenging to predict. Performance trends varied across cities, highlighting temporal differences in mobility patterns that impact downstream tasks. Across all models, Gemini 2.0 Flash consistently achieved the highest accuracy, demonstrating robust zero-shot capabilities across cities and time.
[…]
Model Bandung Beijing Istanbul Jakarta KC Melbourne Moscow NY Palembang PJ SP Shanghai Sydney Tangerang Tokyo
[…]
XGBoost (BoW) 0.620 0.577 0.600 0.635 0.745 0.695 0.670 0.570 0.625 0.580 0.645 0.600 0.600 0.615 0.565
XGBoost (TF-IDF) 0.660 0.615 0.610 0.640 0.735 0.665 0.650 0.565 0.595 0.580 0.685 0.500 0.585 0.620 0.560
[…]
GPT-4.1 Mini 0.585 0.673 0.615 0.600 0.690 0.585 0.745 0.595 0.605 0.575 0.700 0.565 0.515 0.620 0.550
[…]
4.3. Spatiotemporal Classification and Reasoning
This benchmark assesses whether LLMs can be leveraged for spatiotemporal trajectory classification by providing them with contextual information about a user’s behavior. The task evaluates the model’s ability to capture variations in travel patterns across different cities, given the sequence of POI check-ins as input, and without any additional fine-tuning. Through this setup, we aim to understand how effectively LLMs can reason over spatiotemporal and behavioral cues in user trajectories. Appendix F provides details on problem formulation, prompts, and LLM parameters.
4.3.1. Experimental Setup
This task involves trajectory property classification, specifically predicting whether the final check-in occurs on a weekday or a weekend. Each trajectory is converted into a textual prompt incorporating spatial (city), temporal (check-in time-of-day), and categorical contexts (POI category). Adapting the prompt design from LLM-Mob (Wang et al., 2023c), we instructed the LLM to first reason before making a prediction. This approach allows us to evaluate both classification accuracy and the spatiotemporal reasoning capabilities of LLMs, in line with recent work on spatiotemporal reasoning using LLMs (Quan et al., 2025). Whereas prior approaches rely on models that encode trajectories (Nayak and Pandit, 2023), our method directly leverages the LLM’s ability to process contextual information in natural language. We evaluated four closed-source LLM APIs: Gemini 2.0 Flash (Team and et al., 2024a), GPT-4o Mini, GPT-4.1 Mini, and GPT-5 Nano (OpenAI and et al., 2024a, b), and used Accuracy as our primary metric.
4.3.2. Baseline: Bag-of-POI Categories
As a baseline, we also evaluated classical machine learning classifiers: Logistic Regression, Random Forest, and XGBoost (Chen and Guestrin, 2016). For feature extraction, instead of textual prompting, we propose to simply transform each trajectory into a semantic sequence of POI categories. These sequences are then vectorized using either Bag-of-Words (BoW) or TF-IDF, treating POI categories as ”words” and historical user trajectories as a corpus. Unlike the LLMs, these classical models are trained in a fully supervised manner and are very lightweight.
4.3.3. Main Results
As shown in Table 7, Logistic Regression with TF-IDF achieves the highest mean accuracy (0.658) across the 15 cities, outperforming modern LLMs. The best-performing LLM (Gemini 2 Flash) achieves a mean accuracy of 0.645, performing slightly below classical ML baselines on average. This suggests that simple frequency-based representations of categorical features remain highly competitive for this task. Interestingly, the GPT series generally performs worse, even for models that are more recent than Gemini 2 Flash. For instance, GPT-5 Nano records the lowest mean accuracy, despite being designed to support more advanced reasoning capabilities. Overall, these results highlight current limitations of LLMs in capturing spatiotemporal patterns from trajectory data alone. The stronger performance of classical baselines further indicates that, in a zero-shot setting, LLMs still struggle to reason over spatiotemporal patterns as effectively as traditional statistical methods capture categorical regularities.
5. Conclusion and Limitations
Conclusion
In this paper, we presented Massive-STEPS to address longstanding limitations in POI trajectory modeling, particularly the reliance on older, geographically saturated, and non-reproducible check-in datasets. Massive-STEPS offers a large-scale, semantically enriched POI check-in dataset spanning 15 cities across diverse regions and two time periods, supporting single-city, cross-city, and longitudinal studies. The dataset includes rich semantic information about a POI check-in record, including its name, address, category, and geographic coordinates. We also provide comprehensive benchmarking results for supervised POI recommendation, zero-shot POI recommendation, and spatiotemporal classification, demonstrating the dataset’s versatility across models and tasks. Across these tasks, GNNs performed best for supervised POI recommendation, zero-shot LLMs were competitive in some cities, and classical ML models remained strong for spatiotemporal classification. By releasing our dataset and benchmark pipeline publicly, we aim to advance open, reproducible, and geographically inclusive research in POI modeling and human mobility studies.
Limitations
First, Massive-STEPS is derived from Semantic Trails, thus inheriting its biases, which may propagate through downstream tasks. Additionally, the dataset is sparse in several cities, which can impact model training quality and limit cross-city generalization. Second, Massive-STEPS focuses solely on trajectories and POI metadata, without including user demographic or social information due to privacy considerations. This restricts its applicability for personalized or socially-aware POI recommendation tasks. Third, while our benchmarking covers a wide range of models and cities to emphasize replicability and geographic breadth, we did not perform extensive hyperparameter tuning, which may affect the peak performance of the models. Finally, although Massive-STEPS does not reflect present-day mobility patterns, it was designed to provide an alternative to older datasets and to help bridge the gap toward newer, open, and extensible POI benchmarks.
Acknowledgements.
We would like to thank the support of the National Computational Infrastructure and the ARC Center of Excellence for Automated Decision Making and Society (CE200100005). Computational facilities were provided by the School of Computer Science and Engineering at UNSW Sydney through the Wolfpack computational cluster. Additionally, we express our gratitude to the NVIDIA Academic Grant Program for providing access to A100 GPUs on Saturn Cloud. We further thank the Google Cloud Research Credits program for API access to Gemini models.
Appendix A Existing POI Recommendation Datasets
To examine the trend of the usage of POI recommendation datasets, we filtered the comprehensive survey by Zhang et al. (2025) to extract studies that explicitly mention the cities used in their experiments. The resulting distribution is summarized in Table 8, which shows a strong concentration of studies focused on New York and Tokyo. Additionally, Fig. 7 visualizes the same data, highlighting the uneven distribution of city choices across studies. We also include information on the LBSN platforms used, revealing that Foursquare remains the predominant data source in the field. These findings underscore the need for broader, more inclusive datasets that support evaluation across a wider range of global cities.
Table 8. Overview of POI Recommendation Studies by City and LBSN Platform. This table is adapted from Table IV in the survey by Zhang et al. (2025) and presents a filtered list of POI recommendation studies that explicitly mention city names and their associated LBSN platforms.
Study Cities LBSN Dataset Reproducibility Issue
SSTPMF (Davtalab and Alesheikh, 2021) New York, Tokyo Foursquare, Gowalla Gowalla city boundaries not reproducible.
ST-LSTM (Zhao et al., 2018) California, Singapore Brightkite, Foursquare, Gowalla FSQ city boundaries not reproducible. Brightkite and Gowalla not grouped into cities.
LSMA (Wang et al., 2022) New York, San Francisco, Tokyo Foursquare, Weeplaces Weeplaces no longer available.
DLAN (Wu et al., 2024) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
TLR-M (Halder et al., 2021) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
GETNext (Yang et al., 2022c) New York, Tokyo, California Foursquare, Gowalla Only provides preprocessed NYC, missing TKY and CA.
CARAN (Hossain et al., 2022) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
JANICP (Zhong et al., 2022) New York, Tokyo Foursquare, Weeplaces Weeplaces no longer available.
Li et al. (Li et al., 2022a) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
AMACF (Yang et al., 2022a) New York, Tokyo Foursquare, Weeplaces Weeplaces no longer available.
CHA (Zang et al., 2021) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
HAT (Wu et al., 2023) Beijing, Shanghai Yelp, others Used private datasets.
STAR-HiT (Xie and Chen, 2023) New York Foursquare, Gowalla Gowalla not grouped into cities.
CAFPR (Halder et al., 2023) Tokyo, California, Budapest, Melbourne Foursquare Uses POI themepark dataset. POI metadata (lat./lon., category) is missing.
TGAT (Jiang and Wu, 2023) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
MobGT (Xu et al., 2023) New York Foursquare, Gowalla Used private datasets.
POIBERT (Ho and Lim, 2022) Budapest, Delhi, Edinburgh, Glasgow, Osaka, Perth, Toronto Flickr Used private datasets.
AutoMTN (Qin et al., 2022) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
CCDSA (Wang et al., 2023b) New York, Tokyo, San Francisco Foursquare, Weeplaces Weeplaces no longer available.
TDGCN (Cao et al., 2023) Tokyo, California Foursquare, Gowalla, Weeplaces Weeplaces no longer available.
BayMAN (Xia et al., 2023) New York Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
ROTAN (Feng et al., 2024b) New York, Tokyo, California Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
TrajMoE (Han et al., 2025a) Atlanta, Chicago, Seattle, Washington, New York, Los Angeles Unspecified Used private datasets.
UniMove (Han et al., 2025b) Lhasa, Nanchang, Shanghai Unspecified Used private datasets.
STGCN (Han et al., 2020) Boston, Chicago, London Gowalla, others Gowalla city boundaries not reproducible. Used private datasets.
ADQ-GNN (Wang et al., 2021b) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
HS-GAT (Zhang and Ma, 2024) Boston, Chicago, London Yelp, others Yelp not grouped into cities.
HKGNN (Zhang et al., 2024b) New York, Jakarta, Kuala Lumpur, São Paulo Foursquare FSQ city boundaries not reproducible.
S2GRec (Li et al., 2022b) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
GSBPL (Wang et al., 2023a) New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
LSPSL (Jiang et al., 2023) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
SCL (Chen and Zhu, 2025) Florence, Rome, Pisa, Edinburgh, Toronto Flickr Preprocessed Flickr dataset no longer available.
LLM-Move (Feng et al., 2024a) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
LLM4POI (Li et al., 2024) New York, Tokyo, California Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
Refine-POI (Li et al., 2025) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
GNPR-SID (Wang et al., 2025) New York, Tokyo, California Foursquare, Gowalla No issue, uses FSQ-NYC/TKY and Gowalla-CA.
QT-Mob (Chen et al., 2025) New York, Singapore Foursquare, Private Telco Only FSQ-NYC is publicly available.
DiffPOI (Qin et al., 2023) Singapore, New York, Tokyo Foursquare, Gowalla Gowalla not grouped into cities.
DSDRec (Wang et al., 2024) New York, Tokyo Foursquare No issue, uses FSQ-NYC/TKY.
Diff-DGMN (Zuo and Zhang, 2024) Istanbul, Jakarta, São Paulo, New York, Los Angeles Foursquare FSQ city boundaries not reproducible.
Appendix B Data Visualization
We present several visualizations highlighting Massive-STEPS’ scale and diversity to complement our dataset description.
In Fig. 9, we show the top 10 most frequent POI categories for each city. The distribution reflects the local culture and lifestyle across different urban areas. For example, Beijing and Shanghai have a high number of Chinese restaurants, while Melbourne and Sydney show a strong presence of cafés. In Tokyo, convenience stores and ramen shops dominate. These patterns illustrate the diversity of local culture and user interests. Fig. 10 illustrates the temporal shift in the distribution of the top 10 most visited POI categories across each city, comparing the percentage of visits to each category between the 2012-2013 and 2017-2018 periods.
Figure 8. World map highlighting the cities included in the Massive-STEPS dataset.
World map highlighting the cities included in the Massive-STEPS dataset
Table 9. Fields available in the Massive-STEPS dataset, including user, POI, geographic/spatial, and temporal details, along with example data for each field.
Field Description Example
trail_id Numeric identifier of trajectory 2013_2866
user_id Numeric identifier of user 90
venue_id Numeric identifier of POI venue 185
latitude Latitude of POI venue -33.87301862604473
longitude Longitude of POI venue 151.20668402700997
name POI name Sydney Town Hall
address Street address of POI venue 483 George St
venue_category POI category name City Hall
venue_category_id Foursquare Category ID 4bf58dd8d48988d129941735
venue_category_id_code Numeric identifier of POI category 72
venue_city Administrative region name Sydney
venue_city_latitude Latitude of administrative region -33.86785
venue_city_longitude Longitude of administrative region 151.20732
venue_country Country code AU
timestamp Check-in timestamp 2012-04-22 08:20:00
Figure 9. Top 10 most frequent POI categories in each city, highlighting local cultural and urban preferences.
Top 10 most frequent POI categories in each city, highlighting local cultural and urban preferences.
Figure 10. Top 10 most visited POI categories in each city across two time periods, illustrating temporal shifts in user visitation patterns.
Top 10 most visited POI categories in each city across two time periods, illustrating temporal shifts in user visitation patterns.
Appendix C POI Recommendation: Task Details
We adopt the conventional problem formulation used in prior POI recommendation studies (Zhang et al., 2025; Yang et al., 2022c; Yan et al., 2023), which defines the task as learning user preferences and routines from historical check-ins to recommend future POIs.
C.1. Problem Formulation
Let $\mathcal{U}=\{u_{1},u_{2},\dots,u_{M}\}$ denote the set of users, $\mathcal{P}=\{p_{1},p_{2},\dots,p_{N}\}$ the set of Points of Interest (POIs), and $\mathcal{T}=\{t_{1},t_{2},\dots,t_{K}\}$ the set of timestamps, where $M,N,K\in\mathbb{N}$.
POI Definition
Each POI $p\in\mathcal{P}$ is represented as a tuple:
$p=\langle\phi,\lambda,\kappa,\alpha,\beta,\gamma\rangle,$
where:
  • $\phi$ and $\lambda$ are the latitude and longitude,
  • $\kappa$ is the POI category (e.g., restaurant, park),
  • $\alpha$ is the unique POI identifier,
  • $\beta$ is the textual address, and
  • $\gamma$ is the POI name.
Check-in Definition
A check-in is a tuple $c=\langle u,p,t\rangle\in\mathcal{U}\times\mathcal{P}\times\mathcal{T}$, indicating that user $u$ visited POI $p$ at timestamp $t$.
Trajectory Definition
A trajectory for user $u$ is defined as a temporally ordered sequence of check-ins within a fixed time interval $\delta\tau=8$ hours. Each trajectory $T^{i}_{u}(t)$ up to timestamp $t$ is defined as:
$T^{i}_{u}(t)=\{(p_{1},t_{1}),(p_{2},t_{2}),\dots,(p_{k},t_{k})\}$
such that $t_{1}<t_{2}<\dots<t_{k}=t$ and $t_{k}-t_{k-1}\leq\delta\tau$. Given a set of historical trajectories
$\mathcal{T}_{u}=\{T^{1}_{u},T^{2}_{u},\dots,T^{L}_{u}\}$
for user $u$, where $L$ is the number of such historical trajectories, the goal is to recommend the POIs that $u$ is most likely to visit next after the current contextual trajectory $T^{\prime}_{u}(t)$.
POI Recommendation Task Definition
Given a current contextual trajectory $T^{\prime}_{u}(t)$ of user $u$ up to time $t$, along with their historical trajectories $\mathcal{T}_{u}$, the task of next POI recommendation is to rank all candidate POIs $p_{i}\in\mathcal{P}$ according to the model’s predicted probability that user $u$ will visit each POI next.
Formally, the model learns a ranking function:
$f:(T^{\prime}_{u}(t),\mathcal{T}_{u})\rightarrow\{\hat{y}_{i}\}_{i=1}^{|\mathcal{P}|}$
where $\hat{y}_{i}$ denotes the predicted likelihood that user $u$ will visit POI $p_{i}$ next. Based on these scores, a ranked list of POIs is returned as recommendations.
This formulation enables POI recommendation, where the goal is to suggest a set of likely POIs that a user may visit next, based on their historical check-ins and inferred preferences. Our evaluation metrics, Acc@k and NDCG@k, assess whether the ground-truth POI appears among the top-$k$ ranked candidates, reflecting the quality of the recommended set. In particular, Acc@1 captures the stricter task of immediate next POI prediction, measuring whether the top-ranked POI matches the user’s actual next visit.
C.2. Models
For thoroughness, we evaluated the following models as baselines:
  • FPMC (Rendle et al., 2010): A classical baseline that combines first-order Markov chains with matrix factorization to model personalized next-location predictions.
  • RNN (Wang et al., 2021a), LSTPM (Sun et al., 2020), DeepMove (Feng et al., 2018), Flashback (Yang et al., 2021), Graph-Flashback (Rao et al., 2022), and LoTNext (Xu et al., 2024): Recurrent neural networks designed to capture sequential dependencies, with varying mechanisms to incorporate spatio-temporal context.
  • GETNext (Yang et al., 2022c) and STHGCN (Yan et al., 2023): Transformer-based graph neural networks to model social, spatial, and temporal dependencies.
  • UniMove (Han et al., 2025b): Trajectory foundation model based on a Transformer decoder architecture with Mixture of Experts (MoE) layers.
C.3. Experiment and Implementation Details
For training and evaluation, we used the LibCity [^4] https://github.com/libcity/bigscity-libcity-datasets/ library (Wang et al., 2021a), which provides implementations of classical baselines including FPMC (Rendle et al., 2010), RNN (Wang et al., 2021a), LSTPM (Sun et al., 2020), and DeepMove (Feng et al., 2018). The training hyperparameters are listed in Table 10 and, unless otherwise noted, follow the default configurations provided by LibCity.
For Flashback [^5] https://github.com/exascaleinfolab/flashback_code/ (Yang et al., 2021), Graph-Flashback [^6] https://github.com/kevin-xuan/Graph-Flashback (Rao et al., 2022), LoTNext [^7] https://github.com/Yukayo/LoTNext (Xu et al., 2024), GETNext [^8] https://github.com/songyangme/GETNext (Yang et al., 2022c), and STHGCN [^9] https://github.com/alipay/Spatio-Temporal-Hypergraph-Model (Yan et al., 2023), we adapted the source code released by the respective authors. Due to variations in dataset sizes and training costs across cities, we applied different hyperparameters for some cities, as detailed in Table 11.
For UniMove [^10] https://github.com/tsinghua-fib-lab/unimove (Han et al., 2025b), we modified their original source code for Massive-STEPS. For location features, we used Schema.org’s 162 list of categories as a categorical feature and the administrative region as the grid area for POI category distribution. Hyperparameters are listed in Table 10 and, unless otherwise noted, follow the default values.
Table 10. Hyperparameters for Markov-based methods, recurrent networks, and UniMove.
Hyperparameter FPMC RNN LSTPM DeepMove UniMove
Batch Size 20 20 20 20 4
Learning Rate 5e-4 1e-3 1e-4 1e-3 3e-4
Max Epoch 1 30 40 30 50
Location Embedding Size 64 500 500 500 {256, 128}
Hidden Embedding Size N/A 500 500 500 512
Dropout N/A 0.3 0.8 0.5 N/A
Table 11. Hyperparameters for Transformer-based graph neural networks.
Model Cities BS LR Epochs
Flashback, Beijing 10 1e-2 100
Graph-Flashback, Bandung, Istanbul, Jakarta, KC, Melbourne, Moscow, NY 100 1e-2 100
LoTNext Palembang, PJ, SP, Shanghai, Sydney, Tangerang, Tokyo 100 1e-2 100
GETNext Beijing, Melbourne, Moscow, Palembang, Shanghai, Sydney, Tokyo 16 1e-3 200
Bandung, Istanbul, KC, NY, PJ, SP, Tangerang 16 1e-4 20
Jakarta 16 5e-5 20
STHGCN Beijing, Melbourne, Palembang, Shanghai, Sydney, Tokyo 16 1e-4 20
Bandung, Istanbul, Jakarta, KC, Moscow, NY, PJ, SP, Tangerang 64 1e-4 20
All modified code implementations are available as submodules in our main dataset repository. Experiments were conducted using NVIDIA L4, L40S, and H100 GPUs.
C.4. Supplementary Results
We report the full results of our supervised POI recommendation baselines in Table 12, 13 and 14, using three evaluation metrics: Acc@1, Acc@5, and NDCG@5.
Table 12. Performance of supervised POI recommendation baselines across 5 cities: Bandung, Beijing, Istanbul, Jakarta, Kuwait City. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model Bandung Beijing Istanbul Jakarta Kuwait City
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.048 0.118 0.083 0.000 0.021 0.009 0.026 0.074 0.050 0.029 0.085 0.058 0.021 0.089 0.054
RNN 0.062 0.135 0.099 0.085 0.183 0.134 0.077 0.178 0.130 0.049 0.115 0.083 0.087 0.203 0.146
LSTPM 0.110 0.241 0.179 0.127 0.211 0.169 0.142 0.286 0.217 0.099 0.210 0.157 0.180 0.362 0.275
DeepMove 0.107 0.232 0.172 0.106 0.261 0.190 0.150 0.298 0.228 0.103 0.212 0.160 0.179 0.360 0.274
Flashback 0.103 0.202 0.155 0.278 0.435 0.359 0.137 0.257 0.200 0.113 0.214 0.166 0.150 0.298 0.227
Graph-Flashback 0.091 0.197 0.146 0.304 0.426 0.373 0.137 0.261 0.202 0.072 0.167 0.121 0.115 0.245 0.181
LoTNext 0.076 0.169 0.125 0.243 0.391 0.320 0.105 0.227 0.168 0.044 0.110 0.078 0.132 0.280 0.209
GETNext 0.179 0.306 0.247 0.433 0.527 0.486 0.146 0.268 0.210 0.155 0.257 0.209 0.175 0.322 0.251
STHGCN 0.219 0.375 0.302 0.453 0.640 0.552 0.241 0.385 0.318 0.197 0.334 0.270 0.225 0.394 0.314
UniMove 0.007 0.060 0.033 0.036 0.205 0.128 0.015 0.061 0.038 0.004 0.036 0.020 0.023 0.120 0.073
Table 13. Performance of supervised POI recommendation baselines across 5 cities: Melbourne, Moscow, New York, Palembang, Petaling Jaya. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model Melbourne Moscow New York Palembang Petaling Jaya
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.062 0.147 0.107 0.059 0.129 0.094 0.032 0.090 0.061 0.102 0.169 0.136 0.026 0.084 0.057
RNN 0.059 0.105 0.083 0.075 0.164 0.122 0.061 0.119 0.092 0.049 0.121 0.085 0.064 0.148 0.107
LSTPM 0.091 0.204 0.150 0.151 0.300 0.229 0.099 0.206 0.155 0.114 0.230 0.175 0.099 0.222 0.163
DeepMove 0.083 0.179 0.134 0.143 0.283 0.217 0.097 0.195 0.149 0.084 0.191 0.139 0.112 0.234 0.175
Flashback 0.079 0.146 0.115 0.121 0.234 0.180 0.086 0.156 0.123 0.135 0.256 0.198 0.072 0.162 0.119
Graph-Flashback 0.093 0.188 0.143 0.125 0.243 0.188 0.059 0.125 0.093 0.148 0.264 0.208 0.072 0.164 0.119
LoTNext 0.085 0.180 0.133 0.089 0.196 0.145 0.061 0.133 0.099 0.140 0.271 0.208 0.055 0.135 0.097
GETNext 0.100 0.250 0.179 0.175 0.335 0.260 0.134 0.263 0.202 0.158 0.313 0.239 0.139 0.254 0.200
STHGCN 0.168 0.318 0.247 0.223 0.382 0.308 0.146 0.259 0.207 0.246 0.427 0.341 0.174 0.301 0.241
UniMove 0.008 0.066 0.037 0.009 0.051 0.030 0.004 0.028 0.016 0.009 0.060 0.035 0.008 0.058 0.034
Table 14. Performance of supervised POI recommendation baselines across 5 cities: São Paulo, Shanghai, Sydney, Tangerang, Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Model São Paulo Shanghai Sydney Tangerang Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
FPMC 0.030 0.079 0.055 0.084 0.154 0.120 0.075 0.180 0.131 0.104 0.220 0.166 0.176 0.291 0.239
RNN 0.097 0.191 0.147 0.055 0.120 0.090 0.080 0.164 0.125 0.087 0.179 0.135 0.133 0.254 0.197
LSTPM 0.158 0.319 0.243 0.099 0.195 0.149 0.141 0.265 0.206 0.154 0.309 0.237 0.225 0.394 0.315
DeepMove 0.160 0.310 0.240 0.085 0.168 0.128 0.129 0.240 0.188 0.145 0.285 0.219 0.201 0.362 0.288
Flashback 0.159 0.280 0.222 0.096 0.162 0.132 0.127 0.216 0.174 0.185 0.330 0.262 0.170 0.298 0.237
Graph-Flashback 0.096 0.211 0.156 0.111 0.187 0.151 0.135 0.234 0.188 0.200 0.373 0.292 0.180 0.329 0.259
LoTNext 0.105 0.222 0.166 0.093 0.179 0.139 0.123 0.225 0.176 0.126 0.249 0.190 0.191 0.315 0.258
GETNext 0.202 0.360 0.286 0.115 0.230 0.177 0.181 0.347 0.266 0.224 0.372 0.302 0.180 0.361 0.275
STHGCN 0.250 0.425 0.344 0.193 0.329 0.264 0.227 0.378 0.307 0.293 0.492 0.400 0.250 0.432 0.350
UniMove 0.002 0.018 0.009 0.000 0.055 0.029 0.015 0.102 0.059 0.001 0.055 0.029 0.032 0.109 0.072
Appendix D Analyzing Urban Features and POI Recommendation Performance
POI trajectory studies that evaluate models across multiple city-level datasets often include analyses to assess how well their methods generalize across different urban contexts. It is well understood that POI recommendation accuracy metrics (e.g., Acc@k, NDCG@k) can vary substantially between cities and can be interpreted as a proxy for how easy or difficult a city is to model. The assumption is that higher performance reflects more predictable or structured mobility patterns. This viewpoint is consistent with prior work highlighting the role of cultural and urban-specific factors in shaping mobility behaviors (Yang et al., 2015; Sun et al., 2024). For example, Gowalla-CA (Cho et al., 2011; Yuan et al., 2013) often yields lower accuracy compared to FSQ-NYC and FSQ-TKY (Yang et al., 2014), suggesting that some cities may be inherently harder to model.
Several studies have proposed hypotheses connecting specific urban features to modeling difficulty. Yang et al. (2022c) hypothesized that cities with fewer check-ins and higher spatial sparsity of POIs are harder to model. Yan et al. (2023) suggested that a larger number of user trajectories improves predictive accuracy by providing richer collaborative signals, whose architecture is designed to leverage. Li et al. (2024) proposed that cities with a greater variety of POI categories are easier to model due to LLMs’ contextual reasoning capabilities, whereas cities covering a broader geographic area tend to be more difficult to model. In the zero-shot POI recommendation setting, Feng et al. (2025) reported two key findings: (1) geospatial biases inherent in LLMs can hinder prediction quality across cities, and (2) LLMs are influenced by city-specific mobility patterns.
Building on these insights, we used Massive-STEPS to explore how urban features affect POI recommendations. Its diverse set of 15 cities allows for a comprehensive analysis across different cultural and urban contexts. We analyzed the correlation between urban features and model accuracy, and based on the results, proposed a new hypothesis that contrasts previous findings in the literature.
Prior studies (Yang et al., 2022c; Yan et al., 2023; Li et al., 2024) have suggested several features as potential explanatory variables, including:
  • Number of unique check-ins,
  • Number of unique trajectories,
  • Number of unique POI categories,
  • Geographical area (larger areas are assumed to be harder to model), and
  • POI density or spatial sparsity (i.e., unique POIs per unit area).
We also propose several additional features for consideration:
  • Number of unique POIs,
  • Check-in density (unique check-ins per area),
  • Trajectory density (unique trajectories per area), and
  • Category entropy, our proposed feature capturing category diversity.
Category entropy, based on Shannon entropy, measures how evenly POI categories are distributed in a city. A higher entropy suggests that check-ins are spread across a wide variety of categories, while a lower entropy indicates a concentration in fewer types. The formula for Shannon entropy is:
(1) $H=-\sum_{i=1}^{N}p_{i}\log(p_{i})$
where $p_{i}$ is the proportion of venues in category $i$, and $N$ is the total number of POI categories. The proportion $p_{i}$ is defined as:
(2) $p_{i}=\frac{c_{i}}{\sum_{j=1}^{N}c_{j}}$
where $c_{i}$ is the count of venues in category $i$. In other words, $p_{i}$ represents the fraction of all venues that belong to category $i$.
Moreover, previous studies have primarily focused on only three datasets: FSQ-NYC, FSQ-TKY, and Gowalla-CA. In contrast, Massive-STEPS provides broader coverage across 15 cities, enabling a more comprehensive and robust analysis. To examine the relationship between urban features and model performance, we averaged the three evaluation metrics across nine supervised baselines for each city and computed the Spearman correlation with each candidate feature. Fig. 11 presents the correlations between all nine features and the average performance metric.
Among all features, category entropy shows the strongest correlation with model performance, with a Spearman correlation of r=0.646r=-0.646 ($p<0.05$). This suggests that cities with more evenly distributed POI categories tend to yield lower prediction accuracy. Intuitively, when no single category dominates (a city has roughly equal proportions of restaurants, cafes, homes, and other POIs), it becomes more difficult for models to predict a user’s next destination. In these cases, user behavior is more varied and less predictable. On the other hand, cities with more skewed category distributions (e.g., mostly food places or mostly residential areas) tend to have more consistent patterns of movement, making them easier for models to learn and predict. Interestingly, our finding contradicts the hypothesis proposed by LLM4POI (Li et al., 2024), which suggests that FSQ-NYC is easier to model than Gowalla-CA due to the former’s vast number of POI categories, which were supposed to provide richer contextual signals for the model. Our results indicate that it is not the number of categories that matters, but rather how these categories are distributed.
Figure 11. Spearman correlation between nine candidate urban features and the mean score of POI recommendation models across 15 cities.
Spearman correlation between nine candidate urban features and the mean score of POI recommendation models across 15 cities.
Appendix E Zero-shot POI Recommendation: Task Details
E.1. Problem Formulation
The zero-shot POI recommendation task follows the same problem formulation as its supervised counterpart (see Section C.1). The key difference is that in this setting, the model parameters remain frozen and the models are pre-trained, rather than trained from randomly initialized weights.
E.2. Methods
We evaluated three LLM-based prompting methods:
  • LLM-Mob (Wang et al., 2023c): One of the earliest methods to use LLMs for next POI prediction, prompting LLMs with both historical and current (contextual) trajectories.
  • LLM-ZS (Beneduce et al., 2024): A simplified version of LLM-Mob that retains the use of historical and contextual trajectories but simplifies its prompt design.
  • LLM-Move (Feng et al., 2024a): Extends previous prompting methods by introducing a RAG-like approach, retrieving nearby POIs as candidates, and ranking them by geographic distance to the user’s most recent visit.
E.3. Experiment and Implementation Details
Preprocessing
We adopted the AgentMove [^11] https://github.com/tsinghua-fib-lab/agentmove/ library (Feng et al., 2025), which provides implementations of three LLM methods: LLM-Mob (Wang et al., 2023c), LLM-ZS (Beneduce et al., 2024), and LLM-Move (Feng et al., 2024a). The preprocessing steps used by AgentMove are as follows.
First, we selected 200 random users from the test set and sampled one random trajectory for each user. This trajectory serves as the context stays, representing the current trajectory to be predicted. The historical stays are composed of the most recent 15 trajectories from the same user, drawn from the training set. Each check-in is described by four attributes: the hour (in 12-hour format), the day of the week, the POI ID, and the POI category name.
Second, the LLMs are set to return outputs in JSON format, generating the top 5 predicted POI IDs along with an explanation of their reasoning. Following the AgentMove setup and to ensure replicability, we set the generation parameters as follows: a temperature of 0.0, a maximum output length of 1000 tokens, and an input context window capped at 2000 tokens. Following its default configuration, LLM-Move’s candidate set is set to the 100 nearest POIs.
Prompting
Prompt templates for each method, LLM-Mob, LLM-ZS, and LLM-Move, are presented in Listing 1, 2, and 3, respectively.
1Your task is to predict a users next location based on his/her activity pattern.
2You will be provided with <history> which is a list containing this users historical stays, then <context> which provide contextual information
3about where and when this user has been to recently. Stays in both <history> and <context> are in chronological order.
4Each stay takes on such form as (start_time, day_of_week, duration, place_id). The detailed explanation of each element is as follows:
5start_time: the start time of the stay in 12h clock format.
6day_of_week: indicating the day of the week.
7duration: an integer indicating the duration (in minute) of each stay. Note that this will be None in the <target_stay> introduced later.
8place_id: an integer representing the unique place ID, which indicates where the stay is.
9
10Then you need to do next location prediction on <target_stay> which is the prediction target with unknown place ID denoted as <next_place_id> and
11unknown duration denoted as None, while temporal information is provided.
12
13Please infer what the <next_place_id> might be (please output the 10 most likely places which are ranked in descending order in terms of probability), considering the following aspects:
141. the activity pattern of this user that you learned from <history>, e.g., repeated visits to certain places during certain times;
152. the context stays in <context>, which provide more recent activities of this user;
163. the temporal information (i.e., start_time and day_of_week) of target stay, which is important because peoples activity varies during different time (e.g., nighttime versus daytime)
17and on different days (e.g., weekday versus weekend).
18
19Please organize your answer in a JSON object containing following keys:
20"prediction" (the ID of the five most probable places in descending order of probability) and "reason" (a concise explanation that supports your prediction). Do not include line breaks in your output.
21
22The data are as follows:
23<history>: {historical_stays}
24<context>: {context_stays}
25<target_stay>: {target_time, target_day_of_week}
Listing 1: Prompt for LLM-Mob
1Your task is to predict <next_place_id> in <target_stay>, a location with an unknown ID, while temporal data is available.
2
3Predict <next_place_id> by considering:
41. The users activity trends gleaned from <historical_stays> and the current activities from <context_stays>.
52. Temporal details (start_time and day_of_week) of the target stay, crucial for understanding activity variations.
6
7Present your answer in a JSON object with:
8"prediction" (IDs of the five most probable places, ranked by probability) and "reason" (a concise justification for your prediction).
9
10The data:
11<historical_stays>: {historical_stays}
12<context_stays>: {context_stays}
13<target_stay>: {target_time, target_day_of_week}
Listing 2: Prompt for LLM-ZS
1<long-term check-ins> [Format: (POIID, Category)]: {historical_stays}
2<recent check-ins> [Format: (POIID, Category)]: {context_stays}
3<candidate set> [Format: (POIID, Distance, Category)]: {candidates}
4Your task is to recommend a users next point-of-interest (POI) from <candidate set> based on his/her trajectory information.
5The trajectory information is made of a sequence of the users <long-term check-ins> and a sequence of the users <recent check-ins> in chronological order.
6Now I explain the elements in the format. "POIID" refers to the unique id of the POI, "Distance" indicates the distance (kilometers) between the user and the POI, and "Category" shows the semantic information of the POI.
7
8Requirements:
91. Consider the long-term check-ins to extract users long-term preferences since people tend to revisit their frequent visits.
102. Consider the recent check-ins to extract users current perferences.
113. Consider the "Distance" since people tend to visit nearby pois.
124. Consider which "Category" the user would go next for long-term check-ins indicates sequential transitions the user prefer.
13
14Please organize your answer in a JSON object containing following keys:
15"prediction" (10 distinct POIIDs of the ten most probable places in <candidate set> in descending order of probability), and "reason" (a concise explanation that supports your recommendation according to the requirements). Do not include line breaks in your output.
Listing 3: Prompt for LLM-Move
Models and Implementations
We use the following LLMs in our experiments:
All open-source models are quantized using AWQ (Lin et al., 2024) and served via vLLM (Kwon et al., 2023). Inference of open-source models was conducted on NVIDIA A100 GPUs. We accessed Gemini via the official API. All modified code implementations are publicly available in our main dataset repository.
E.4. Supplementary Results
We provide the full results of our zero-shot POI recommendation results in Table 15, 16, and 17, providing three metrics: Acc@1, Acc@5, and NDCG@5. Additionally, Table 18, 19, and 20 present zero-shot performance across two time periods (2012-2013 and 2017-2018) using LLM-Move.
Table 15. Performance of zero-shot POI recommendation baselines across 5 cities: Bandung, Beijing, Istanbul, Jakarta, Kuwait City. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Method Model Bandung Beijing Istanbul Jakarta Kuwait City
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.105 0.170 0.139 0.115 0.308 0.226 0.080 0.225 0.160 0.100 0.245 0.174 0.095 0.270 0.185
Qwen 2.5 7B 0.060 0.155 0.111 0.058 0.385 0.218 0.035 0.240 0.148 0.105 0.245 0.179 0.080 0.220 0.155
Llama 3.1 8B 0.010 0.100 0.055 0.000 0.000 0.000 0.020 0.110 0.065 0.055 0.150 0.104 0.030 0.100 0.066
Gemma 2 9B 0.070 0.175 0.126 0.115 0.288 0.206 0.075 0.200 0.146 0.105 0.240 0.178 0.080 0.210 0.150
LLM-ZS Gemini 2 Flash 0.095 0.195 0.147 0.058 0.385 0.246 0.090 0.235 0.166 0.110 0.250 0.188 0.080 0.245 0.167
Qwen 2.5 7B 0.055 0.185 0.126 0.038 0.404 0.237 0.040 0.235 0.141 0.065 0.250 0.161 0.050 0.220 0.140
Llama 3.1 8B 0.045 0.210 0.131 0.077 0.346 0.221 0.040 0.225 0.137 0.045 0.200 0.126 0.060 0.210 0.137
Gemma 2 9B 0.065 0.185 0.130 0.096 0.308 0.217 0.045 0.225 0.141 0.105 0.250 0.180 0.070 0.230 0.153
LLM-Move Gemini 2 Flash 0.225 0.350 0.289 0.096 0.346 0.218 0.205 0.385 0.289 0.295 0.405 0.350 0.220 0.380 0.295
Qwen 2.5 7B 0.100 0.155 0.128 0.192 0.346 0.280 0.175 0.270 0.226 0.115 0.225 0.169 0.160 0.285 0.227
Llama 3.1 8B 0.030 0.035 0.033 0.058 0.135 0.100 0.015 0.055 0.036 0.015 0.025 0.021 0.010 0.035 0.023
Gemma 2 9B 0.175 0.245 0.213 0.096 0.365 0.229 0.100 0.200 0.155 0.235 0.290 0.266 0.120 0.275 0.202
Table 16. Performance of zero-shot POI recommendation baselines across 5 cities: Melbourne, Moscow, New York, Palembang, Petaling Jaya. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Method Model Melbourne Moscow New York Palembang Petaling Jaya
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.060 0.150 0.108 0.130 0.245 0.187 0.095 0.175 0.136 0.135 0.275 0.208 0.090 0.220 0.160
Qwen 2.5 7B 0.030 0.130 0.083 0.090 0.270 0.185 0.070 0.185 0.131 0.075 0.205 0.143 0.030 0.195 0.116
Llama 3.1 8B 0.010 0.065 0.040 0.030 0.100 0.068 0.025 0.090 0.061 0.005 0.040 0.025 0.010 0.090 0.050
Gemma 2 9B 0.055 0.150 0.108 0.100 0.240 0.176 0.070 0.175 0.124 0.095 0.240 0.171 0.055 0.185 0.122
LLM-ZS Gemini 2 Flash 0.065 0.160 0.115 0.125 0.300 0.217 0.080 0.170 0.129 0.130 0.260 0.196 0.110 0.210 0.164
Qwen 2.5 7B 0.040 0.155 0.100 0.080 0.260 0.176 0.050 0.180 0.116 0.050 0.215 0.135 0.045 0.175 0.111
Llama 3.1 8B 0.040 0.155 0.101 0.080 0.270 0.183 0.055 0.160 0.111 0.070 0.240 0.154 0.030 0.205 0.123
Gemma 2 9B 0.050 0.140 0.100 0.080 0.290 0.194 0.075 0.185 0.129 0.060 0.235 0.150 0.065 0.185 0.126
LLM-Move Gemini 2 Flash 0.225 0.325 0.275 0.220 0.400 0.316 0.235 0.415 0.325 0.260 0.385 0.329 0.210 0.335 0.273
Qwen 2.5 7B 0.110 0.220 0.165 0.230 0.310 0.274 0.120 0.255 0.188 0.130 0.195 0.163 0.135 0.175 0.155
Llama 3.1 8B 0.040 0.195 0.123 0.005 0.065 0.031 0.035 0.130 0.084 0.010 0.015 0.013 0.040 0.060 0.049
Gemma 2 9B 0.115 0.275 0.199 0.110 0.245 0.185 0.115 0.245 0.183 0.210 0.270 0.240 0.175 0.235 0.208
Table 17. Performance of zero-shot POI recommendation baselines across 5 cities: São Paulo, Shanghai, Sydney, Tangerang, Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Method Model São Paulo Shanghai Sydney Tangerang Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
LLM-Mob Gemini 2 Flash 0.130 0.305 0.223 0.055 0.160 0.111 0.060 0.160 0.112 0.155 0.285 0.225 0.140 0.320 0.238
Qwen 2.5 7B 0.090 0.290 0.188 0.040 0.170 0.108 0.035 0.145 0.091 0.095 0.285 0.196 0.110 0.350 0.243
Llama 3.1 8B 0.030 0.165 0.098 0.005 0.020 0.013 0.020 0.085 0.053 0.020 0.120 0.073 0.005 0.045 0.025
Gemma 2 9B 0.085 0.230 0.162 0.050 0.150 0.104 0.030 0.130 0.086 0.145 0.270 0.209 0.145 0.345 0.255
LLM-ZS Gemini 2 Flash 0.150 0.315 0.235 0.065 0.160 0.113 0.060 0.155 0.111 0.145 0.310 0.234 0.160 0.380 0.278
Qwen 2.5 7B 0.095 0.290 0.198 0.045 0.155 0.103 0.045 0.170 0.109 0.100 0.315 0.215 0.120 0.365 0.257
Llama 3.1 8B 0.030 0.280 0.159 0.060 0.165 0.116 0.040 0.185 0.110 0.080 0.255 0.173 0.110 0.415 0.269
Gemma 2 9B 0.075 0.300 0.192 0.050 0.165 0.112 0.045 0.155 0.103 0.100 0.330 0.227 0.110 0.395 0.263
LLM-Move Gemini 2 Flash 0.285 0.415 0.350 0.170 0.270 0.221 0.230 0.420 0.331 0.200 0.340 0.274 0.250 0.470 0.368
Qwen 2.5 7B 0.155 0.235 0.199 0.095 0.165 0.133 0.125 0.280 0.205 0.175 0.280 0.229 0.250 0.360 0.312
Llama 3.1 8B 0.045 0.045 0.045 0.020 0.040 0.030 0.055 0.220 0.141 0.000 0.005 0.003 0.030 0.060 0.046
Gemma 2 9B 0.195 0.300 0.252 0.105 0.150 0.128 0.125 0.370 0.254 0.125 0.250 0.193 0.130 0.305 0.225
Table 18. Performance of zero-shot POI recommendation using LLM-Move across two time periods and 5 cities: Bandung, Beijing, Istanbul, Jakarta, and Kuwait City. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Time Period Model Bandung Beijing Istanbul Jakarta Kuwait City
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
2012-2013 Gemini 2 Flash 0.227 0.351 0.290 0.102 0.367 0.232 0.212 0.384 0.290 0.295 0.409 0.352 0.423 0.500 0.453
Qwen 2.5 7B 0.098 0.155 0.126 0.204 0.367 0.298 0.192 0.295 0.247 0.114 0.223 0.167 0.269 0.423 0.357
Llama 3.1 8B 0.031 0.036 0.034 0.041 0.122 0.086 0.007 0.027 0.017 0.010 0.021 0.016 0.000 0.000 0.000
Gemma 2 9B 0.180 0.247 0.217 0.102 0.388 0.244 0.116 0.199 0.159 0.228 0.285 0.260 0.308 0.385 0.342
2017-2018 Gemini 2 Flash 0.167 0.333 0.272 0.000 0.000 0.000 0.185 0.389 0.284 0.286 0.286 0.286 0.190 0.362 0.271
Qwen 2.5 7B 0.167 0.167 0.167 0.000 0.000 0.000 0.130 0.204 0.168 0.143 0.286 0.233 0.144 0.264 0.208
Llama 3.1 8B 0.000 0.000 0.000 0.333 0.333 0.333 0.037 0.130 0.088 0.143 0.143 0.143 0.011 0.040 0.026
Gemma 2 9B 0.000 0.167 0.083 0.000 0.000 0.000 0.056 0.204 0.142 0.429 0.429 0.429 0.092 0.259 0.181
Table 19. Performance of zero-shot POI recommendation using LLM-Move across two time periods and 5 cities: Melbourne, Moscow, New York, Palembang, Petaling Jaya. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Time Period Model Melbourne Moscow New York Palembang Petaling Jaya
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
2012-2013 Gemini 2 Flash 0.226 0.329 0.279 0.218 0.401 0.316 0.240 0.403 0.320 0.256 0.385 0.327 0.199 0.348 0.274
Qwen 2.5 7B 0.116 0.232 0.175 0.234 0.310 0.275 0.130 0.240 0.190 0.128 0.195 0.162 0.142 0.184 0.164
Llama 3.1 8B 0.039 0.200 0.125 0.005 0.066 0.032 0.032 0.117 0.074 0.010 0.015 0.013 0.014 0.043 0.028
Gemma 2 9B 0.097 0.271 0.189 0.112 0.249 0.188 0.130 0.260 0.198 0.215 0.272 0.244 0.199 0.262 0.234
2017-2018 Gemini 2 Flash 0.222 0.311 0.261 0.333 0.333 0.333 0.217 0.457 0.342 0.400 0.400 0.400 0.237 0.305 0.273
Qwen 2.5 7B 0.089 0.178 0.132 0.000 0.333 0.167 0.087 0.304 0.183 0.200 0.200 0.200 0.119 0.153 0.134
Llama 3.1 8B 0.044 0.178 0.116 0.000 0.000 0.000 0.043 0.174 0.116 0.000 0.000 0.000 0.102 0.102 0.102
Gemma 2 9B 0.178 0.289 0.232 0.000 0.000 0.000 0.065 0.196 0.133 0.000 0.200 0.086 0.119 0.169 0.144
Table 20. Performance of zero-shot POI recommendation using LLM-Move across two time periods and 5 cities: São Paulo, Shanghai, Sydney, Tangerang, Tokyo. We report three metrics: Acc@1 (A@1), Acc@5 (A@5), and NDCG@5 (N@5).
Time Period Model São Paulo Shanghai Sydney Tangerang Tokyo
A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5 A@1 A@5 N@5
2012-2013 Gemini 2 Flash 0.298 0.440 0.369 0.192 0.288 0.242 0.256 0.462 0.367 0.197 0.338 0.272 N/A N/A N/A
Qwen 2.5 7B 0.173 0.250 0.215 0.109 0.186 0.151 0.122 0.288 0.209 0.172 0.278 0.226 N/A N/A N/A
Llama 3.1 8B 0.048 0.048 0.048 0.006 0.026 0.016 0.064 0.224 0.148 0.000 0.005 0.003 N/A N/A N/A
Gemma 2 9B 0.202 0.315 0.264 0.109 0.160 0.136 0.122 0.378 0.257 0.126 0.253 0.195 N/A N/A N/A
2017-2018 Gemini 2 Flash 0.219 0.281 0.251 0.091 0.205 0.147 0.136 0.273 0.204 0.500 0.500 0.500 0.250 0.470 0.368
Qwen 2.5 7B 0.063 0.156 0.115 0.045 0.091 0.070 0.136 0.250 0.192 0.500 0.500 0.500 0.250 0.360 0.312
Llama 3.1 8B 0.031 0.031 0.031 0.068 0.091 0.077 0.023 0.205 0.117 0.000 0.000 0.000 0.030 0.060 0.046
Gemma 2 9B 0.156 0.219 0.189 0.091 0.114 0.101 0.136 0.341 0.246 0.000 0.000 0.000 0.130 0.305 0.225
Appendix F Spatiotemporal Classification and Reasoning: Task Details
F.1. Problem Formulation
Borrowing the notation used in Section C, we formulate this task as follows. Given a current contextual trajectory $T^{\prime}_{u}(t)$ of user $u$ up to time $t$, the goal of spatiotemporal trajectory classification is to predict a property $y$ of the trajectory. In this study, we focus on weekday/weekend classification, where $y\in\{\text{weekday},\text{weekend}\}$.
Formally, the LLM serves as a classification function:
$f:T^{\prime}_{u}(t)\rightarrow\hat{y}$
where $\hat{y}$ denotes the predicted class label for the trajectory. The model is evaluated based on its accuracy in correctly classifying trajectories according to this property.
F.2. Experiment and Implementation Details
Preprocessing
We borrowed the experimental setup of AgentMove, similar to our zero-shot POI recommendation procedure in Section E.3. That is, we selected 200 random users from the test set and sampled one random trajectory for each user. This trajectory is then included in the test set. Each check-in is described by four attributes: the hour (in 12-hour format), the day of the week, the POI ID, and the POI category name.
LLMs are set to return outputs in a structured/JSON format, predicting whether the trajectory ended on a weekday or a weekend, along with an explanation of their reasoning. To ensure replicability, Gemini and GPT-4 models are set with the following generation parameters: a temperature of 0.0, a maximum output length of 1000 tokens, and an input context window capped at 2000 tokens. Due to API requirements, GPT-5 Nano uses the following generation parameters: a fixed temperature of 1.0, a maximum output length of 4096 tokens, low verbosity, and medium reasoning effort.
Prompt
Prompt template for spatiotemporal weekday-weekend classification is shown in Listing 4.
1A trajectory is a sequence of check-ins, each represented as (start_time, poi_category). The detailed explanation of each element is as follows:
2start_time: the start time of the check-in in 12h clock format.
3poi_category: the category of the point of interest (POI) visited during the check-in
4
5The trajectory is as follows: {[check-in time-of-day, POI category] for check-in in trajectory}
6
7Your task is to classify whether the last check-in occurs on a weekday or a weekend.
8Consider the temporal information (i.e., start_time) of the trajectory, which is important because peoples activity varies during different time (e.g., nighttime versus daytime).
9Consider the POI categories, which can provide insights into the users activity patterns.
10Also consider the city context, as different cities may have different cultural and social norms that influence activity patterns. The city is: {city}.
11
12Please organize your answer in a JSON object containing following keys:
13"prediction" ("weekday" or "weekend") and "reason" (a concise explanation that supports your prediction).
14Do not include line breaks in your output.
Listing 4: Prompt for Weekday vs. Weekend Classification
Models and Implementations
We use the following LLMs in our experiments:
  • Gemini 2.0 Flash (gemini-2.0-flash),
  • GPT-4o Mini (gpt-4o-mini),
  • GPT-4.1 Mini (gpt-4.1-mini),
  • GPT-5 Nano (gpt-5-nano).
We accessed Gemini and GPT models via the official API. All modified code implementations are publicly available in our main dataset repository.
Appendix G License and Data Usage
Our work does not involve the collection of new data. Instead, we derive our resulting dataset by combining and aligning two publicly available datasets, both of which are distributed under permissive licenses. We did not scrape data from the internet or use proprietary APIs to construct this dataset.
We accessed the Semantic Trails Dataset (Monti et al., 2018) via Figshare: https://doi.org/10.6084/m9.figshare.7429076.v2. The dataset is licensed under the Creative Commons CC0 1.0 license (https://creativecommons.org/publicdomain/zero/1.0/), which allows unrestricted copying, modification, and redistribution for any purpose, including commercial use, without requiring permission.
We accessed the Foursquare Open Source Places dataset via Hugging Face: https://huggingface.co/datasets/foursquare/fsq-os-places. Foursquare Open Source Places is licensed under the Apache License, Version 2.0. More details are available in Foursquare’s documentation: https://docs.foursquare.com/data-products/docs/access-fsq-os-places:
Copyright 2024 Foursquare Labs, Inc. All rights reserved. Licensed under the Apache License, Version 2.0 (the ”License”); you may not use this file except in compliance with the License. You may obtain a copy of the License at:  http://www.apache.org/licenses/LICENSE-2.0 Unless required by applicable law or agreed to in writing, software distributed under the License is distributed on an ”AS IS” BASIS, WITHOUT WARRANTIES OR CONDITIONS OF ANY KIND, either express or implied. See the License for the specific language governing permissions and limitations under the License.
We release our Massive-STEPS dataset under the same Apache Version 2.0 License, and have included Foursquare Open Source Places’ license in our hosted dataset’s README file.